### v1

# Taylor’s hypothesis and its impact on flux measurements in a forest clearcut

Abstract
Taylor’s hypothesis is the backbone to convert observations done over time to spatial information of the flow while carrying out turbulence measurements on a micrometeorological tower. To address its validity over a highly heterogeneous forest clearcut surface, we utilize an extensive Distributed Temperature Sensing (DTS) and Eddy Covariance (EC) datasets. The DTS measured space-time correlation curves of temperature fluctuations are used to compute the bulk convective speeds ($U_{C}$) of temperature structures in buoyant conditions at a height of $\approx$ 3.1 m above the clearing. These convective speeds are compared with the mean wind speed ($U$) and turbulent intensities of streamwise velocities ($\sigma_{u}\textfractionsolidus U$) obtained from the EC system at the middle of the clearcut. Depending on if $U$ is parallel or perpendicular to the forest edge, the relationships between $U_{C}\textfractionsolidus U$ and $\sigma_{u}\textfractionsolidus U$ are significantly different. However, irrespective of the wind direction, the convective speeds of temperature structures at inertial subrange scales behave in a power-law fashion with increasing wavenumbers. The exponent of this power-law ($-0.8$) differs from a homogeneous atmospheric surface layer flow ($-1/3$), thereby pointing towards the effects of heterogeneity. The scale-dependent convective speeds non-linearly transform the temporal frequencies to streamwise wavenumbers, which, eventually, impacts the properties of turbulence (co) spectra. More importantly, this non-linear transformation yields a critical frequency limit, beyond which the eddy length scales derived from frequencies are smaller than the physical dimension of the sonic anemometers, and therefore, cannot be faithfully resolved. This critical limit questions the EC flux estimates beyond this frequency.
[…]
The two major components of the Earth’s surface energy budget are sensible and latent heat fluxes. Sans spatial information, the accuracy of these flux measurements is tied to Taylor’s hypothesis, which states that the space-time conversion of any turbulence dataset can be accomplished through the local mean wind speed. We test Taylor’s hypothesis over a highly heterogeneous forest clearcut surface by using fiber-optic cables that acquired air temperatures at high spatial and temporal resolutions. From this data, we estimate the speeds of turbulent eddies at each spatial scale of the flow and derive a critical frequency limit beyond which the space-time conversion is highly uncertain. This critical limit questions the turbulent flux estimates beyond this frequency.
1 Introduction
The quantification of vertical turbulent transport of heat, matter (e.g. moisture, $\rm CO_{2}$) and momentum are important for understanding the land-atmosphere interaction, energy budget of the Earth’s surface, and carbon and water exchanges between the ecosystem and atmosphere. Under the assumptions of homogeneity and stationary conditions, these vertical transports can be interpreted as surface fluxes, which eventually serve as boundary conditions for the Earth System Models. Given such various applications, the micrometeorological towers at different locations across the globe routinely monitor these fluxes under the aegis of various programs such as FLUXNET, AmeriFlux, NEON, and ICOS (Baldocchi et al., 2001; Novick et al., 2018; Metzger et al., 2019; Heiskanen et al., 2022). Although the turbulent eddies facilitating the vertical transport exist physically in space, spatial information of atmospheric turbulent flows are rarely available and thus the instruments mounted on micrometeorological towers register the turbulent variables over time (Hilland and Christen, 2024). To circumvent this issue, Taylor’s frozen turbulence hypothesis is used to convert the temporal information to spatial ones. This hypothesis states that the turbulent structures of any spatial dimension move past the measurement location with a convection velocity in a frozen state, i.e. their characteristics change slowly compared to the speed at which the mean flow transports them along. In the absence of spatial information, it remains experimentally challenging to verify the accuracy of the Taylor’s hypothesis (TH) under various land-surface conditions.
Despite so, a handful of previous studies attempted to verify TH either through theoretical formulations or by conducting novel field experiments that provided spatial information. From the theoretical side, the early study of Wyngaard and Clifford (1977) used a model developed by Lumley (1965) to investigate the effects of TH on the turbulence statistics at smaller scales of the flow. In particular, Wyngaard and Clifford (1977) provided theoretical formulations that accounted for the deviation from the assumption of constant convective speeds across all scales of the flow. Later on, Tong (1996) experimentally verified their results by using two spatially separated ultrasonic anemometers. Except a few experimental set-ups, spatially explicit sonic anemometer measurements are not readily available (Han et al., 2019). Some studies deployed thermocouples to test TH but their spatial resolution was rather limited (Deardorff and Willis, 1982; Everard et al., 2021). However, with the advent of novel instrumentation, spatial data of turbulent variables can be obtained at high enough resolution either from remote sensing or through fiber optic cables. In that regard, Higgins et al. (2012) used a Raman Lidar to investigate the convective speeds of temperature and moisture structures over a landscape surrounding a water body. On the other hand, Cheng et al. (2017) used distributed temperature sensing (DTS) measurements over a grassland to verify the hypothesis for different atmospheric stability conditions.
The DTS measurement technique uses fiber optic cables and relies on Raman scattering principle to convert the back-scattered energy from the emitted light pulses to temperature values at specific locations in space (Peltola et al., 2021). By laying out these cables either horizontally or vertically, it is therefore possible to obtain temperature data at a fine enough spatial (of the order of tens of cm) and temporal resolutions (of the order of seconds). Cheng et al. (2017) used this spatial temperature data from a horizontal configuration to evaluate scale-dependent convective speeds of turbulent structures. For their experimental dataset over a flat grass surface, they found that at smaller scales of the flow, the turbulent structures moved at speeds that scaled with the streamwise wavenumbers ($\kappa$) as ${(\kappa)}^{-1/3}$. To reach this conclusion, Cheng et al. (2017) adopted the technique of Del Alamo and Jiménez (2009), where differences in Fourier phases of spatial temperature signals at two subsequent instant of time were used to estimate the wavenumber-dependent convection speeds. Cheng et al. (2017) also provided theoretical support to this $-1/3$ exponent by borrowing formulations from the literature of homogeneous and isotropic turbulence. They hypothesized that these scale-dependent convective speeds could enhance the turbulent fluxes by 10-20% at smaller scales of the flow, which, eventually, could solve the crisis of surface energy imbalance.
In recent times, DTS measurements were used over a flat and homogeneous Namib desert surface to verify TH (Hilland and Christen, 2024). Hilland and Christen (2024) found that the TH was almost never satisfied in atmospheric surface layer flows under various stability conditions (ranging from buoyancy- to shear-dominated flows), and the measurement height was an important parameter in that regard. Contrary to Cheng et al. (2017), Hilland and Christen (2024) used a correlation-cube methodology to estimate the scale-dependent convective speeds. In their method, the definition of scales is not related to streamwise wavenumbers in a Fourier domain. In spite of this, Hilland and Christen (2024) found that the convective speeds of smaller-scale turbulent structures were indeed lower than the bulk convective speeds of larger-scale structures. Be that as it may, Hilland and Christen (2024) did not report any power-law variations in their estimates of scale-dependent convective speeds.
All these studies described above were conducted over homogeneous atmospheric surface layer flows. However, in reality, one mostly encounter strong heterogeneity in land-surface conditions, which presents a significant challenge to the atmospheric community (Bou-Zeid et al., 2020). One of such examples is a forest clearcut. Clearcutting is a widely utilized forest management practice to collect timber and in some cases also to restrict the spread of wildfires by limiting the access to the fuel (Ortega et al., 2024). From a fluid dynamical perspective, clearcutting introduces additional complexities to the air flows. For instance, when the air blows directly from the forest canopy sub-layers to clearcut location, it encounters a significant change in surface roughness, which causes the flow to deviate from being in equilibrium with local surface characteristics. The laboratory experiments of Chung and Koseff (2021) over an artificial heterogeneous canopy showed that these deviations are often caused by the canopy-scale eddies extending their footprints on the clearcut surface depending on the distance from the canopy edge. On the other hand, the field-experimental results of Detto et al. (2008) provided evidence that the flow near the forest edge shares several attributes with a backward-facing step flow in which a recirculation zone dominates the turbulence statistics.
From a climate science standpoint, the forests are strong sources and sinks of water vapor and carbon dioxide, respectively, thereby influencing the water- and carbon-cycles between the atmosphere and the ecosystem. Forest clearcutting clearly impacts these exchange processes and the climatic impact of such forest management practices can be observed with the eddy covariance (EC) technique (Tikkasalo et al., 2025). However, due to the aforementioned complexities, the turbulent behavior of the air flow in forest clearcuts pose challenges while interpreting the observed flux values. One of such aspects is regarding the applicability of TH. For instance, if the convective speeds of turbulent structures are indeed scale-dependent then it can have serious consequences on the EC flux measurements as hypothesized by Cheng et al. (2017). Furthermore, the relation derived by Cheng et al. (2017) to describe the scale-dependent convective speeds was under the assumption of homogeneous and isotropic turbulence. A recent study by Chowdhuri and Banerjee (2024) revealed that the turbulence is hardly isotropic at smaller scales of the flow corresponding to the heights closer to the surface. Moreover, it is uncertain whether the assumption of homogeneity would remain valid in highly heterogeneous conditions, such as in a forest clearcut. To address these pertinent issues, our objectives in this study are primarily threefold:
  1. 1.
    Estimate the bulk convective speeds of turbulent structures from the DTS data and use that information to validate TH.
  2. 2.
    Investigate how the convective speeds vary with increasing wavenumbers.
  3. 3.
    Quantify the impact of scale-dependent convective speeds on turbulence statistics and discuss its implications for EC flux measurements and surface energy balance.
To tackle these objectives, we employ an extensive dataset encompassing both DTS and EC measurements, collected during a five-month field campaign in a forest clearcut. In our analysis, we specifically consider two scenarios depending on whether the direction of the mean wind was perpendicular or parallel to the forest edge. In the latter scenario, when the air blows parallel to the forest edge, it does not encounter sharp changes in surface roughness, and therefore, is expected to behave differently than a forest-edge flow. To estimate the convective speeds of the turbulent structures, the space-time correlation curves are constructed from the DTS temperature measurements. This method has been adopted previously for laboratory (Deardorff and Willis, 1982), large eddy simulation (Su et al., 2000), and field-experimental (Han and Zhang, 2022) datasets. In addition to that, we introduce a novel concept of conditional space-time correlation curves to separately estimate the convective speeds of positive and negative temperature structures. More importantly, we develop a comprehensive framework that accounts for deviations from TH on turbulence spectra and cospectra. This framework, in general, can be applied to situations other than a forest clearcut. The rest of this article is organized in three different sections. In Sect. 2 we describe our measurement set up and methodologies being adopted to verify TH. In Sect. 3, we present and discuss our results, and in Sect. 4 future research directions are outlined.
2 Data and Methodology
2.1 Site description
The measurement campaign took place between May 17 and September 10 2024 at Ränskälänkorpi study site (6111 N, 2516 E, 144 m above sea level), which is located on a peatland that was drained before 1960’s for forestry purposes (Laurila et al., 2021). A Norway spruce (Picea abies) stand growing on the site was clear-felled during late winter - spring of 2021 creating a 6.4 ha clearcut area (see Fig. 1) and Norway spruce seedlings were planted after the clearcutting. Few isolated groupings of living and dead trees were left untouched on the clearcut area and logging debris (e.g. branch piles) were left on site. In general, the terrain of the clearcut is flat, but its surface exhibits prominent small-scale undulations due to e.g. drainage ditches and banks. After clearcutting, pioneer species, such as fireweed (Chamaenerion angustifolium), downy birch (Betula pubescens) and raspberry (Rubus idaeus), have spread on the area and created a varying vegetation cover on the clearing (see Fig. 1a). During the study period, the height of these plants varied approximately between 0.5 m and 1.5 m, depending on location and timing of the growing season. In summary, the clearcut area is a complex mosaic of isolated trees, pioneer species, and logging debris creating a rough surface for the flow.
The clearcut is surrounded by a forest dominated by Norway spruce (Picea abies) with height between 20 m and 25 m depending on the location (see Fig. 1b). Plant area index of the forest in SW direction of the clearing is approximately xx m2 m-2. The forest edges create abrupt changes in canopy height and hence in surface roughness.
Figure 1: (a) Image and a (b) map of the study site. The image was taken towards South-West on June 28 2024 at the location of the white triangle in (b). Different DTS cable sections are highlighted with red lines. In (b) the black contours show the topography with 2-m increments based on the National Land Survey of Finland Topographic Database (data retrieved: 10/2023). The canopy height map was derived from a drone flight conducted during September 27 2022 and complemented with canopy height model of Finnish Forest Center (data retrieved: 12/2024). Note that the canopy height shown in the map does not fully reflect the conditions at the clearing during the study period due to the rapidly developing vegetation cover.
2.2 Measurement setup
The spatial and temporal variability of air temperature above the clearcut was measured with the distributed temperature sensing (DTS) technique Thomas and Selker (2021). The DTS instrument (Ultima-S, 5 km variant, Silixa Ltd., Hertfordshire, UK) was housed in a temperature controlled cabin in the middle of the forest clearcut (see Fig. 1b). The measurements were conducted using double-ended configuration and signals from both ends of the fiber-optic cable (forward and reverse signals) were stored for later processing (Sect. 2.3). The DTS instrument provided observations with 1.2 s temporal and 12.7 cm spatial resolution along the attached cable. Note that in double-ended configuration both ends are measured sequentially and hence after processing the temperature observations were available with 2.3 s resolution.
Two calibration locations were created next to the DTS instrument and both ends of the cable were guided through them providing altogether four calibration sections in the data. The calibration sections were at minimum 10 meters long. Cold calibration point was created by digging the cable to 0.7 m depth in the peat soil. This depth was below the soil water level and hence the temperatures at that depth did not vary rapidly. Typically, the temperatures were below the air temperature and varied between 2.9 ${}^{\circ}C$ and 10.7 ${}^{\circ}C$ during the campaign. Warm calibration point was created with 50-l calibration bath filled with water and it was located in the temperature controlled cabin. Water heater was installed on June 28 and it was set to keep the water at 30 C. Prior to that the water temperature followed indoor air temperature in the temperature controlled cabin. Temperatures at the calibration locations were monitored with reference thermometers (PT-100) supplied with the DTS instrument.
A white, thin (outer diameter 0.9 mm) aramid reinforced 50 $\mu$m multimode fiber-optic cable (AFL Telecommunications, Duncan, SC 29334, US) was utilized in this study. The cable was arranged in six sections and along each section the temperature values were doubly measured. This was because the measurements were made with a single continuous cable traversing back and forth the different sections (Fig. 1). Sections 1 to 5 were aligned with the forest edge located in South-West direction of the clearcut and the section lengths were between 97 m and 103 m. The cable sections 1, 2, 3, 4 and 5 were approximately at a 20 m, 47 m, 75 m, 115 m and 137 m distance from the forest edge in SW direction, respectively. The sixth section was perpendicular to the other cable sections and to the forest edge and its length was 164 m, out of which 27 m was within the forest. Therefore, the total cable length was 1478 m.
The cable was fastened to wooden support structures and kept at a constant height throughout the measurement domain. The cable locations were determined with a handheld GNSS device (Geo 7X, Trimble Inc., USA). This height was around 3.1 m above the soil, which is similar to the height of the location where eddy covariance (EC) measurements were made (see Fig. 1b). However, due to the heterogeneity in the surface conditions (see Sect. 2.1), local DTS measurement height varied and was approximately between 2 to 4 m above the soil and between 0.5 to 2.5 m above the local roughness elements (e.g. vegetation). Moreover, occasionally the cable was also surrounded by the few isolated trees left at the clearcut.
To quantify the turbulent fluxes over such a complex surface, the EC equipment consisted of a 3-D sonic anemometer (uSonic-3 Cage MP, METEK GmbH, Germany) and an Infrared gas analyzer (LI-7200RS, LI-COR Biosciences, NE, USA). These together measured the three wind velocity components, sonic temperature, and carbon dioxide and water vapor mixing ratios at the same height as DTS, i.e. 3.1 m above the soil. The infrared gas analyzer sampled the flow next to the sonic anemometer and all the EC data were logged with 10-Hz frequency. See more details on the EC measurement setup in Tikkasalo et al. (2025).
2.3 DTS data processing
DTS data were post-field calibrated with Python (3.11.9) scripts and the processing followed des Tombe et al. (2020), Van De Giesen et al. (2012) and Peltola et al. (2022) with slight differences. Temperature can be estimated from the observed Stokes ($P_{s}$) and anti-Stokes ($P_{as}$) signals with the following equation (Van De Giesen et al., 2012):
[…]
Noise estimates ($\sigma_{T}$, 1-$\sigma$ variability related to random noise) of $T_{f}$ and $T_{r}$ observations were computed for each 30-min period as,
$\sigma_{T}^{2}(s)=\frac{\overline{\left(T(s,t)-T(s+\Delta s,t)\right)^{2}}}{2},$ (2)
where $\overline{()}$ denotes temporal averaging and $\Delta s$ is the spatial step along the cable (0.127 cm). In this equation, we assume that the variability between consecutive data points in space is dominated by noise. The calibrated temperature $T$ at each $t$ and $s$ was calculated as a weighted average of $T_{f}$ and $T_{r}$ where inverse of noise estimates were used as weights (des Tombe et al., 2020). More importantly, the noise estimates for $T$ were calculated following noise propagation of weighted averaging. The spatial resolution of the data were reduced to 30 cm with spatial aggregation using weighted averaging as above.
After this calibration procedure, we obtained calibrated $T$ at each time step $t$ and location $s$ along the cable. The $T$ data measured along the cable were converted to physical coordinates (ETRS89/TM35FIN coordinate system) using the measured cable locations (Sect. 2.2). Since there were double cabling at each physical location, the double measurements made at each location were further averaged using weighted averaging as above. By utilizing this procedure, we obtained calibrated $T$ data with 2.3 s temporal and 30 cm spatial resolution throughout the measurement campaign at well-defined physical ($x$,$y$,$z$) locations (red lines in Fig. 1) over the study area. These calibrated temperature values were compared against the sonic anemometer measurements at the clearcut showing satisfactory agreement between the two (not shown).
2.4 Convection speed estimation
Figure 2: (a) The space-time contour plot of temperature fluctuations ($T^{\prime}$) is shown from the DTS measurements of a specific section (Section-1) for a particular 30-min period between 10:30-11:00 UTC on 18 May 2024. The spatial locations along the cable section are denoted by $X$ and the temperature fluctuations are obtained by linearly detrending the temperature signals measured along time. The white horizontal strips indicate the regions where the DTS cable was fastened to masts and the temperature values were omitted from those locations. (b) The space-time correlation curves of temperature fluctuations ($R_{TT}(r,\tau)$) are shown for the same 30-min period obtained from Section-1. The temporal lags or leads are denoted as $\tau$ while the colors indicate different values of spatial separations along $X$, $r$. The blue squares denote the peaks of the correlation curves ($\tau_{p}$) for each value of $r$. (c) A scatter plot between $r$ and $\tau_{p}$ is shown. The thick black color denotes the best fit line to compute the convective speed of the temperature structures.
To use the DTS and EC datasets, we focus on buoyancy-dominated conditions when the temperature signals were sufficiently strong. In particular, we selected those 30-min periods when there was no rain, incoming shortwave radiation was larger than 300 $\rm W\ m^{-2}$ and it was not strongly varying within the 30-min periods, the sensible heat flux ($>50\rm\ Wm^{-2}$) and variance of temperature ($>0.1\rm\ K^{2}$) were sufficiently large, air temperature did not exhibit strong trends, and the direction of the mean wind was within ±$\pm 20^{\circ}$ of the DTS cable positions. This wind direction criterion resulted in two different scenarios. For Sections 1 to 5, the mean wind was parallel to the forest edge while for Section 6 it was perpendicular to the edge (see Fig. LABEL:fig:1b). After applying all these conditions, we ended up with a total of 35 hours of DTS and EC data, or in other words, 70 half-hour periods. The DTS dataset was used to estimate the convective speeds of the turbulent structures by using the temperature signal as a tracer. The EC dataset served as a testbed to evaluate the impact of these convective speeds on the routine flux measurements.
Out of all our selected runs, for a specific representative 30-min period (10:30-11:00 UTC, 18 May 2024), we show the space-time contour plot of temperature fluctuations from Section 1 of the DTS cable (Fig. 2a). The temperature values were omitted from those locations where the DTS cable was fastened to wooden structures. Nevertheless, to ensure continuity in the data, those missing values were replaced with linearly-interpolated ones. Thereafter, the temperature fluctuations ($T^{\prime}$) were computed by linearly detrending the data over time separately at each location along the cable sections. The distance along Section 1 is denoted as $X$. To compute the convective speed of the temperature structures, we construct the space-time correlation curves from the DTS dataset. Mathematically, the two-point correlation functions of temperature fluctuations can be defined as,
$R_{TT}(\tau,r)=\frac{\overline{T^{\prime}(X,t)T^{\prime}(X+r,t+\tau)}}{\sigma_{T(X)}\sigma_{T(X+r)}},$ (3)
where $\tau$ are the temporal lags or leads, $r$ denotes the spatial separations along $X$, and $\sigma_{T(X)}$, $\sigma_{T(X+r)}$ are the standard deviations of temporal temperature fluctuations at locations $X$ and $X+r$. For the same data as shown in Fig. 2a, the quantity $R_{TT}(\tau,r)$ are plotted in Fig. 2b. Here, different colors represent the increasing values of $r$ while the blue squares indicate the peaks ($\tau_{p}$) attained by $R_{TT}(\tau,r)$ for a specific $r$. By fitting a straight line to a scatter plot between $r$ and $\tau_{p}$, the bulk convective speed of the temperature structures can be estimated (see the solid black line in Fig. 2c). The goodness of the fit can be adjudged by its $R^{2}$ value, which was well-beyond 0.9 in this example case (Fig. 2c). This linear relationship between $\tau_{p}$ and $r$ can be described as, $r=U_{c}\tau_{p}$, where $U_{C}$ is the bulk convective speed ($\rm m\ s^{-1}$). We restrict the $r$ values within 21 m while making the scatter plot, since for $r>21\rm\ m$, the $R_{TT}(\tau,r)$ values are small and therefore the estimations of $\tau_{p}$ are not statistically robust.
To test the robustness of this method for atmospheric surface layer flows, we repeated this analysis on the Cheng et al. (2017) dataset and the results yielded similar features (see Fig. S1 of the Supplementary Material). However, this particular method does not distinguish between positive and negative structures, or the possible scale-dependence of the convection velocities. In convective flows, a significant asymmetry is observed between the positive and negative temperature structures, which carry the signatures of the buoyant updrafts and downdrafts (Chowdhuri et al., 2020b). This asymmetry is related to how the flow is organized under such conditions. We illustrate this aspect in Fig. 9 of Appendix A. Therefore, it is prudent to investigate whether these two structures move at similar speeds. In the absence of spatial vertical velocity measurements, we associate these updrafts and downdrafts with positive and negative temperature fluctuations, and refer to them as up- and down-plumes, respectively.
Figure 3: The contours of joint probability density function (JPDF) of temperature fluctuations separated by a spatial distance $r$ are shown in subplot (a) for the same period as used in Fig. 2. The contour values are represented by a colorbar while the different quadrants signify how the signs of the temperature fluctuations co-vary. The thick dash-dotted black line represents a straight line with a slope equal to the Pearson correlation coefficient between the two temperature signals. The space-time correlation curves are shown for two different scenarios, by conditioning the temperature fluctuations based on (b) when they are only positive, i.e. relate to the quadrant I in subplot (a) ($R^{+}_{TT}(r,\tau)$) or (c) when they are only negative, i.e. relate to the quadrant III ($R^{-}_{TT}(r,\tau)$). The colors indicate different values of $r$ and the pink and blue squares denote the peaks of the conditional correlation curves ($\tau^{+}_{p}$ and $\tau^{-}_{p}$) for each value of $r$. (d) A scatter plot between $r$ and the peak positions of the conditional correlation curves is shown. The solid pink and blue colors denote the best fit lines to compute the convection speeds of the positive and negative temperature structures.
To carry out that exercise, we first look at the joint probability density functions (JPDFs) of temperature fluctuations at two distinct spatial locations along Section 1. These are denoted as, $T^{\prime}(X+r)$ and $T^{\prime}(X)$ and subsequently normalized by their respective standard deviations. From Fig. 3a it is clear that the JPDFs between $T^{\prime}(X+r)/\sigma_{T_{X+r}}$ and $T^{\prime}(X)/\sigma_{T_{X}}$ is skewed towards the third quadrant (III), where $T^{\prime}(X+r)<0,T^{\prime}(X)<0$. In other words, there exists a bias between the quadrants I ($T^{\prime}(X+r)>0,T^{\prime}(X)>0$) and III, which potentially could cause differences between the convective speeds of positive and negative temperature values. To lend credence to this hypothesis, we construct the conditional space-time correlation curves where $R_{TT}(r,\tau)$ values are computed based on whether $T^{\prime}(X+r)$ and $T^{\prime}(X)$ lie in quadrant I or III of Fig. 3a. This can be expressed as,
$R^{+}_{TT}(\tau,r)=\frac{\overline{T_{\rm I}^{\prime}(X,t)T_{\rm I}^{\prime}(X+r,t+\tau)}}{\sigma_{T_{\rm I}(X)}\sigma_{T_{\rm I}(X+r)}},R^{-}_{TT}(\tau,r)=\frac{\overline{T_{\rm III}^{\prime}(X,t)T_{\rm III}^{\prime}(X+r,t+\tau)}}{\sigma_{T_{\rm III}(X)}\sigma_{T_{\rm III}(X+r)}}.$ (4)
The subscripts I and III in Eq. 4 indicate that the temperature fluctuations are either sampled from the positive or negative quadrants. In Figs. 3b–c, $R^{+}_{TT}(\tau,r)$ and $R^{-}_{TT}(\tau,r)$ curves are shown for different values of $r$ and $\tau$. The pink and blue squares denote the peaks of $R^{+}_{TT}(\tau,r)$ and $R^{-}_{TT}(\tau,r)$ curves at a specific $r$ value. These peak positions are represented as $\tau^{+}_{p}$ and $\tau^{-}_{p}$, respectively. Similar to Fig. 2c, the straight line fits to the scatter plots of $r$ and $\tau^{\pm}_{p}$ (with $R^{2}>0.9$) are used to compute the sign-dependent convection speeds, expressed as, $U^{+}_{c}$ and $U^{-}_{c}$ respectively (Fig. 3d).
The same procedure was repeated for the other sections of the DTS cable to obtain $U_{c}$, $U^{+}_{c}$, and $U^{-}_{c}$. Specific to cable Section 6, we ignored the portion of the cable that resided within the forest. This was done to alleviate any additional complexities associated with sub-canopy scale eddy motions.
2.5 Scale-dependent convection speed
To compute the scale-dependent convection speeds of temperature structures, we follow the same procedure as outlined in Cheng et al. (2017). By borrowing the concept of phase-dependent convection speeds from Del Alamo and Jiménez (2009), Cheng et al. (2017) defined the convection speed at a wavenumber $\kappa$ ($U_{c}(\kappa)$) as,
$U_{c}(\kappa)=\frac{\Delta\varphi}{\kappa\times\Delta t}$ (5)
Here, $\Delta\varphi$ indicates the Fourier phase differences between the spatial temperature signals at time instants $t$ and $t+\Delta t$, $\kappa$ is the spatial wavenumber directly evaluated from the DTS measurements ($\kappa=2\pi/\lambda$, where $\lambda$ is the wavelength), and $\Delta t$ is the temporal lag (2.3 s for our DTS data). Unlike Cheng et al. (2017), we used a wavelet based approach (Torrence and Compo, 1998) to compute the phase differences between the spatial temperature signals at times $t$ and $t+\Delta t$, i.e. $\Delta\varphi$ For our purposes, Morlet was used as the mother wavelet and upon comparison with a Fourier based approach, both yielded similar results (see Fig. S4a in the Supplementary material). Specifically, a wavelet based approach allowed us to estimate the scale-dependent convection speeds separately for the positive and negative temperature structures. To do so, corresponding to a particular wavelet scale, the wavelet coefficients of the spatial temperature signals at a time $t$, were conditionally averaged only over those spatial locations where the temperature fluctuations were positive. An analogous operation was carried out for the negative counterparts and for a time $t+\Delta t$. We, therefore, defined the scale-dependent connection speeds of positive and negative structures as,
$U^{+}_{c}(\kappa)=\frac{\Delta\varphi_{\rm P}}{\kappa\times\Delta t},U^{-}_{c}(\kappa)=\frac{\Delta\varphi_{\rm N}}{\kappa\times\Delta t},$ (6)
In Eq. 6, the phase differences were only evaluated conditionally over the locations where temperature fluctuations were either positive or negative. This is denoted by the subscripts P and N, respectively.
3 Results and discussion
We begin with demonstrating the relationship between the convective speed estimations from the DTS measurements and the mean wind speed from the EC system at the clearcut. Corresponding to the EC data, $u$ and $w$ indicate the streamwise and vertical velocities, while $T,q,c$ indicate temperature, moisture, and $\rm CO_{2}$ concentrations. After employing a double-coordinate rotation, the turbulent fluctuations in these quantities are computed through linear detrending. By using both DTS and EC datasets, the impact of scale-dependent convection speeds on the turbulence statistics are discussed. We end by pointing out some potential limitations of the DTS measurements and how our findings influence in general the flux measurements from the eddy covariance systems.
3.1 Convection speeds of temperature structures
3.1.1 Bulk convection speeds
Figure 4: (a) The scatter plot between the convection speed estimated from the DTS measurements ($U_{C}$) and the mean wind speed from EC system ($U$) at the clearcut. The dash-dotted black line show the 1:1 relationship while the blue dash-dotted line is the best fit line. (b)–(c) The ratios $U_{c}/U$ are plotted against the turbulence intensities $\sigma_{u}/U$ for the cases when the direction of the mean wind was either along or perpendicular to the forest edge. For the perpendicular case, the wind directions are further subdivided when the wind was approaching directly from the forest or from the clearcut. (d) The ratios between the convection speeds of positive ($U^{+}_{c}$) and negative ($U^{-}_{c}$) temperature structures are shown against the skewness of temperature fluctuations obtained from the EC measurements. (e) By combining the EC and DTS measurements, a scatter plot between the length scales of positive and negative temperature structures is shown. Here, $z$ is the measurement height, $\gamma^{+}_{T}$ and $\gamma^{-}_{T}$ are the integral time scales of positive and negative temperature fluctuations while $U^{+}_{c}$ and $U^{-}_{c}$ denote their convection speeds, respectively.
Figure 4a shows the scatter plot between the convective speed of temperature structures ($U_{C}$) estimated from the space-time correlation curves (Fig. 2b–c) and the mean wind speed from the EC system ($U$) at the clear cut. The gray-colored squares denote the Sections 1 to 5, and these correspond to the scenario when the wind was blowing parallel to the forest edge (along-wind). The red-colored squares denote Section 6 and correspond to the scenario when the direction of the mean wind was perpendicular to the forest edge (perpendicular-wind). The black dash-dotted line indicates 1:1 agreement between $U_{C}$ and $U$. For the along-wind cases, it is clear that the $U_{C}$ values are consistently larger than $U$, albeit with some scatter (gray-colored squares in Fig. 4a). Therefore, the temperature structures are advected faster than the local mean wind speed and the relationship between $U_{C}$ and $U$ can be expressed as, $U_{C}=0.77U+1.04$ (see the blue dash-dotted line in Fig. 4a). The same relationship describes the perpendicular-wind cases as well (red-colored squares), but the scatter is significantly increased.
According to the study by Willis and Deardorff (1976), the Taylor’s hypothesis is rendered invalid when the turbulent intensities of the streamwise velocities ($\sigma_{u}/U$, where $\sigma_{u}$ is the standard deviation) exceed 0.5. In Figs. 4b–c, we show the scatter plots between $U_{c}/U$ and $\sigma_{u}/U$. Although the number of samples are limited, it appears that for the along-wind cases (Fig. 4b), $U_{c}/U$ indeed remains close to 1 when $\sigma_{u}/U\leq 0.5$. However, for values other than 0.5, the relationship between $U_{c}/U$ and $\sigma_{u}/U$ is linear with an asymptote of 1.9 as $\sigma_{u}/U>0.8$. All these relationships can be condensed to a singular form,
$U_{c}/U=\mathcal{H}(0.5-\sigma_{u}/U)+(3\sigma_{u}/U-0.5)[\mathcal{H}(\sigma_{u}/U-0.5)-\mathcal{H}(\sigma_{u}/U-0.8)]+1.9\mathcal{H}(\sigma_{u}/U-0.8),$ (7)
where $\mathcal{H}(x)$ is a Heaviside function with $\mathcal{H}(x)=0$ when $x\leq 0$ and 1 otherwise.
For the perpendicular-wind cases, the mean wind directions can be further separated into two categories; when the wind was blowing directly from the forest and when it approached the forest edge from the clearcut. For the first category (forest), it appears that the ratios $U_{c}/U$ were closer to unity despite the changes in $\sigma_{u}/U$ (Fig. 4c). Therefore, corresponding to this flow category, the Taylor’s hypothesis appears to be valid even for $\sigma_{u}/U>0.5$. However, for the latter category (clearcut), the pink squares in Fig. 4c followed Eq. 7 more closely. This difference is not trivial by considering the fact that when the wind blows from the forest, the turbulence at the clearcut location has not fully equilibrated to the local surface conditions. On the other hand, when the wind approaches the forest edge from the clearcut, the situation reverses.
[…]
3.1.2 Conditional convection speeds
Moving on, to see whether the positive and negative temperature structures move at the same speeds, we computed the ratio $U^{+}_{C}/U^{-}_{C}$, where $U^{+}_{C}$ and $U^{-}_{C}$ were estimated from the conditional space-time correlation curves (Figs. 3b–d). From a statistical perspective, the asymmetry between positive and negative temperatures is related to the skewness of temperature fluctuations. As shown in Fig. 9a of Appendix A, for our selected periods, the distributions of temperature fluctuations are positively skewed. In Fig. 4d, the ratios $U^{+}_{C}/U^{-}_{C}$ were plotted against the skewness of temperature fluctuations ($\mathcal{S}(T)$) computed from the EC measurements. This was done to check if the statistical asymmetry could explain any differences between $U^{+}_{C}$ and $U^{-}_{C}$. Note that in Fig. 4d there are fewer data-points corresponding to the perpendicular-wind cases. This is because the peaks ($\tau^{\pm}_{p}$) of the conditional space-time correlation curves did not move nicely with $r$ in order to have robust estimates of $U^{+}_{C}$ and $U^{-}_{C}$. For the along-wind cases, $U^{+}_{C}/U^{-}_{C}$ values were averaged over all the five sections and the spread between them is shown as error-bars with one standard deviation. From Fig. 4d it appears that, for both along- and perpendicular-wind cases, the ratios $U^{+}_{C}/U^{-}_{C}$ remain at a constant value of 0.8 irrespective of temperature skewness.
Since $U^{+}_{C}/U^{-}_{C}<1$, the up-plumes tend to move slower than the down-plumes. In essence, this result suggests that the temporal statistics are based on a biased sample of turbulent flow when compared against sampling over spatial domain, which, however, is not in practical terms possible. In terms of boundary-layer physics, the down-plumes bring colder air from aloft to the layer close to the surface and since these plumes carry the signatures of the large-scale structures, they appear to move faster (Li et al., 2018). We now use this critical information to compute the averaged sizes of up- and down-plumes from the EC measurements. To do that, we first estimate the auto-correlation functions conditioned upon whether the temperatures are positive or negative. Mathematically, these functions can be defined as,
$R^{+}_{TT}(\tau)=\frac{\overline{[T^{\prime}(t)|>0][T^{\prime}(t+\tau)|>0]}}{\sigma^{2}_{T}},$ (8)
where $R^{+}_{TT}(\tau)$ indicate the autocorrelation functions of temperature fluctuations conditioned on their positive values only. $R^{-}_{TT}(\tau)$ is defined in a similar way but for the negative ones. By integrating $R^{+}_{TT}(\tau)$ and $R^{-}_{TT}(\tau)$ up to their first zero crossing, one can estimate the integral time scales of the positive ($\gamma^{+}_{T}$) and negative ($\gamma^{-}_{T}$) temperature structures. The functions $R^{+}_{TT}(\tau)$ and $R^{-}_{TT}(\tau)$ are compared with the original autocorrelation functions ($R_{TT}(\tau)$) in Fig. S3 of the Supplementary Material. By multiplying $\gamma^{+}_{T}$ and $\gamma^{-}_{T}$ with their respective convection speeds ($U^{+}_{C}$ and $U^{-}_{C}$), the physical dimensions of the up- and down-plumes can be calculated.
In Fig. 4e, we show a scatter plot between $(\gamma^{+}_{T}U^{+}_{C})/z$ and $(\gamma^{-}_{T}U^{-}_{C})/z$, where $z$ is the measurement height ($z=3.1$ m). We only show this comparison for the along-wind cases and one can clearly see that the sizes of the down-plumes can be as large as 12 times the measurement height. In a statistical sense, the down-plumes remain nearly 2.5 times larger than the up-plumes (see the blue dash-dotted line in Fig. 4e). This information is important e.g. from the perspective of choosing the domain size for the large eddy simulation models, since the down-plumes would require considerably larger area for their statistics to converge.
3.1.3 Scale-dependent convection speeds
Figure 5: (a) The scale-dependent convection speeds ($U_{c}(\kappa z)$) are plotted against the streamwise wavenumbers ($\kappa$) computed directly from the DTS spatial information. The wavenumbers are normalized by the measurement height ($\kappa z$) while $U_{c}(\kappa z)$ values are scaled by the bulk convection speed $U_{C}$. The gray- and red-colored lines denote the along- and perpendicular-wind cases, respectively. The error bars indicate the spread, computed as one standard deviation from the ensemble mean. The various fits to the scale-dependent convection speeds are shown by the dash-dotted lines (see the legend). (b) Same as in (a), but $U_{c}(\kappa z)$ values are scaled by the convection speed estimated from an elliptical model of space-time correlation curves. (c)–(d) The scale-dependent convection speeds are shown separately for the positive and negative temperature structures.
Hitherto, we have discussed the bulk convective speeds but now we turn our attention towards the scale-dependent convection speeds. In Fig. 5a, we show the ensemble-averaged scale-dependent convection speed estimations from the spatial DTS data of temperature at all the six sections. As before, the Sections 1–5 correspond to along-wind cases while Section 6 represents the perpendicular ones. The convection speeds in Fig. 5a are normalized by the bulk convective speed $U_{C}$ and the wavenumbers ($\kappa$) computed from the spatial DTS separations are scaled with the measurement height ($z$). The choice of $z$ to scale $\kappa$ is in line with Hilland and Christen (2024), who claimed that in the context of Taylor’s hypothesis the measurement height is indeed an important parameter. Ideally, one should use $z-d$, where $d$ is the displacement height, but here we used only $z$ since to estimate $d$ a detailed vertical profile of wind measurements is required and that was not available at our measurement site. This would not have any impact on the results as $d$ only acts to shift the $x$ coordinates horizontally.
Figure 5a depicts that at lower wavenumbers, $\kappa z<1$, $U_{C}(\kappa z)$ values tend towards $U_{C}$, irrespective of the wind directions. On the other hand, at larger wavenumbers, $\kappa z>1$, $U_{C}(\kappa z)/U_{C}$ values decrease in a power-law fashion with the exponent being equal to ${(\kappa z)}^{-0.8}$ (Fig. 5a). This is surprisingly different from the result of Cheng et al. (2017) who reported this exponent to be ${(\kappa z)}^{-1/3}$. We replotted the results of Cheng et al. (2017) and indeed for different atmospheric stability conditions (near-neutral to highly-convective) their data agree more closely with ${(\kappa z)}^{-1/3}$ instead of ${(\kappa z)}^{-0.8}$ (see Fig. S4b in the Supplementary material). However, compared to ours, the data of Cheng et al. (2017) exhibited a significant amount of scatter.
Cheng et al. (2017) theoretically derived this $-1/3$ exponent for atmospheric surface layer flows by exploiting the advection-diffusion equation and a power-law dependence of the diffusion coefficient under the assumption of homogeneous and isotropic turbulence at the inertial subrange scales. For our case, the apparent disagreement with their results could be attributed to two factors. First, we consider highly-heterogeneous flow conditions such as a forest clearcut. Second, the recent results of Chowdhuri and Banerjee (2024) have demonstrated that, for a wide range of both atmospheric and laboratory datasets, significant anisotropy can persist even at smaller scales of the flow. This is true for our EC dataset, as the flow remains anisotropic at the inertial subrange scales when computed through the intermittency-anisotropy framework of Chowdhuri and Banerjee (2024) (not shown).
For atmospheric surface layer and sparse canopy flows, a couple of previous studies have shown that an elliptic model of space-time correlation curves to compute the convection speeds is a better approach as opposed to a linear one adopted here (Han and Zhang, 2022; Everard et al., 2021). We describe this elliptical model in Appendix B. In this model, the convective speeds of the turbulent structures are estimated as $\sqrt{U_{e}^{2}+V_{e}^{2}}$, where $U_{e}$ and $V_{e}$ are convection and sweeping velocities, respectively. In Fig. 5b, $\sqrt{U_{e}^{2}+V_{e}^{2}}$ is used to normalize $U_{C}(\kappa z)$ and the results demonstrate that under this scaling the scale-dependent convection speeds do not exactly approach the bulk convective speeds ($\sqrt{U_{e}^{2}+V_{e}^{2}}$) at lower wavenumbers. Furthermore, the collapse among different sections is poorer as compared to Fig. 5a. Therefore, this questions the applicability of an elliptic model in these heterogeneous flow conditions.
To understand the origin of $-0.8$ exponent in scale-dependent convection speeds, we conditioned them based on the instances when the spatial temperature fluctuations were either positive or negative (see Sect. 2.5). For consistency, both $U^{+}_{C}(\kappa z)$ and $U^{-}_{C}(\kappa z)$ values were scaled with $U_{c}$. From Figs. 5c–d, it is clear that the $-0.8$ exponent is connected to the convective speeds of the down plumes ($U^{-}_{C}(\kappa z)$), contrary to the up ones ($U^{+}_{C}(\kappa z)$). In fact, at lower wavenumbers ($\kappa z<1$), $U^{-}_{C}(\kappa z)$ remain larger than $U^{+}_{C}(\kappa z)$ values and at larger wavenumbers ($\kappa z>1$) the convective speeds of the up plumes decrease as fast as ${(\kappa z)}^{-1.5}$. Therefore, as per the terminology of Cheng et al. (2017), the down-plumes diffuse slower than the up-plumes while carried along with the flow. We next discuss the implications of the scale-dependent convection speeds on the estimations of turbulence spectra and cospectra from EC measurements conducted in a temporal domain.
3.2 Impact on turbulence spectra
3.2.1 Frequency to wavenumber conversion
Under the assumption that the turbulent structures of different scales advect past the measurement location at similar speeds, one could essentially transform the frequencies ($f$) to wavenumbers ($\kappa$) using $\kappa=2\pi f/U$, where $U$ is the mean wind speed estimated from the sonic anemometer. Irrespective of whether one uses $U$ or $U_{C}$ to convert the frequencies to wavenumbers, it is a linear transformation. However, if the convective speeds of turbulent structures are scale-dependent, it introduces additional complexities to the transformation. From our DTS observations, it is clear that $U_{C}(\kappa z)$ is equal to $U_{C}$ at scales $\kappa z<1$, while it varies as ${(\kappa z)}^{-0.8}$ at scales $\kappa z>1$ (see Fig. 5a). One could express this dependence through the following functional form,
$\frac{U_{c}(\kappa z)}{U_{c}}=\frac{1}{{(1+\kappa z)}^{4/5}},$ (9)
which satisfies the aforementioned conditions. When compared against the observations, this functional form fits the data exceptionally well (see the yellow dash-dotted line in Fig. 5a). Now, if this relationship is used to convert the frequencies to wavenumbers, one could write,
$\kappa z=\frac{2\pi fz}{U_{C}(\kappa z)}.$ (10)
By replacing Eq. 9 in Eq. 10, and after some algebraic manipulations, one could rewrite Eq. 10 as,
${(\kappa z)}^{5}=a{(1+\kappa z)}^{4},$ (11)
where $a=(2\pi fz/U_{C})^{5}$. By solving the polynomial in Eq. 11 and restricting its roots to positive values, one could essentially convert the frequencies to wavenumbers by using a scale-dependent convection speed.
Before we discuss the impact of this transformation on turbulence spectra and cospectra, it is important to notice that since Eq. 11 is a non-linear function, the wavenumbers generated from Eq. 11 will not be spaced equally in a linear space. Accordingly, the wavenumber increments ($d\kappa$) themselves would depend on the values of $\kappa$. In fact, by differentiating Eq. 11, one can show,
$d\kappa=\frac{10\pi}{U_{C}}\frac{{(1+\kappa z)}^{9/5}}{5+\kappa z}df.$ (12)
Eq. 12 dictates that $d\kappa=f(\kappa)$ at a fixed $z$, meaning that constant $df$ does not convert to constant $d\kappa$. This creates an issue while estimating the spectral densities ($S_{xx}(\kappa)$) as those are defined as $S_{xx}(\kappa)=E_{xx}(\kappa)/d\kappa$, where $E_{xx}(\kappa)$ are the spectral powers of a turbulent signal $x$ and $d\kappa$ is the wavenumber increment independent of $\kappa$.
Therefore, to solve this paradox, we exponentially bin the scaled wavenumbers ($\kappa z$) generated from Eq. 11 and then compute the bin-averaged spectral powers ($E_{xx}(\kappa z_{\rm bin})$) corresponding to those wavenumber ranges, bounded between $\kappa z_{\rm bin}(i)<\kappa z<\kappa z_{\rm bin}(i)+d\log(\kappa z_{\rm bin})$, where $i$ spans from 1 to $N$ with $N$ being the number of exponential bins being used. Here, we use $N=50$ for our application. To get an equivalent estimate in the linear space, the spectral density at a wavenumber of $\kappa z_{\rm bin}(i)$ is expressed as, $S_{xx}(\kappa z_{\rm bin}(i))=E_{xx}(\kappa z_{\rm bin}(i))/[\kappa z_{\rm bin}(i)d(\log(\kappa z_{\rm bin})]$, where $\kappa z_{\rm bin}(i)$ itself appears in the denominator. This is necessary since $d(\log(\kappa))=d\kappa/\kappa$. The entire procedure described above is a standard one while dealing with empirical estimations of the probability density functions of stochastic variables that are known to display power-law variations (Chowdhuri et al., 2020a). Henceforth, while reporting the spectral results derived from the scale-dependent convection speeds, we omit the subscript from $\kappa z_{\rm bin}$.
3.2.2 DTS temperature spectra
Figure 6: The ensemble averaged temporal and spatial spectra of temperature fluctuations from the DTS measurements are shown for (a)–(e) along- and (f) perpendicular-wind cases. Subplots (a)–(e) show data from cable sections 1 to 5 and subplot (f) shows data from section 6. The spectral amplitudes ($S_{TT}$) are normalized with the variance of temperature ($\sigma^{2}_{T}$) computed from the EC measurements at the clearcut. The error bars indicate the spread, computed as one standard deviation from the ensemble mean. For the temporal spectra, the frequency ($f$) to wavenumber ($\kappa$) conversion is carried out using bulk ($U_{C}$) and scale-dependent ($U_{c}(\kappa z)$) convection speeds, and shown as blue and black colored lines, respectively. For the spatial data, the wavenumbers are directly computed from the available spatial information (shown as red-colored lines). The Kolmogorov $-5/3$ prediction for the spectral slopes is shown for both temporal and spatial measurements.
To test whether this procedure generates meaningful results, we used the DTS dataset of temperature, since that encompasses both temporal and spatial information. First, we computed the spectra of temperature fluctuations in a temporal domain, where the frequencies were converted to wavenumbers ($\kappa=2\pi f/U_{C}$) using the bulk convective speed $U_{C}$ (see the blue squares in Fig. 6). Second, the spectra of temperature fluctuations were computed from the spatial data, where the wavenumbers ($\kappa=2\pi/\lambda$, where $\lambda$ is the wavelength) were estimated directly from the spatial distances (red squares in Fig. 6). Third, the temporal spectra of temperature were replotted using the scale-dependent convective speeds to convert the frequencies to wavenumbers (see Eq. 11) while the spectral densities were calculated using the exponential bin framework (black squares in Fig. 6). In Fig. 6, we show these three comparisons for all the six sections of the DTS cable. As noted before, Sections 1-5 describe the along-wind periods (Figs. 6a–e) while Section 6 describes the perpendicular-wind periods (Fig. 6f). Notwithstanding the sections, the spectra in temporal space agree exceptionally well with the spatial spectra, when the frequencies are converted to wavenumbers using Eq. 11. More importantly, a clear $-5/3$ spectral slope is recovered in the temporal domain under this transformation.
3.2.3 EC spectra and cospectra
Figure 7: The ensemble-averaged (a–b) spectra and (c–d) cospectra of scalars (temperature, moisture, and $\rm CO_{2}$) and velocity ($u$ and $w$) components are shown from the clearcut EC measurements, pertaining to the cases when the wind was blowing parallel to the forest edge (along-wind). In (b) and (d), the conversion of frequency to wavenumber is carried out using the scale-dependent convection speeds, as estimated from the DTS data. (e)–(h) The same information as in (a)–(d), but for the perpendicular-wind cases. The spectral and cospectral slopes of $-1$, $-5/3$, and $-7/3$ are shown as solid lines on the respective panels.
After establishing this, we do the same exercise with the EC measurements. Here, we focus separately on the along-wind (Figs. 7a–d) and perpendicular-wind periods (Figs. 7e–h). The spectral and cospectral densities ($S_{xx}$ and $S_{wx}$, where $x=u,w,T,q,c$) in Fig. 7 are normalized with the respective variances and product of the standard deviations. If the mean wind speed from the sonic anemometer ($U$) is used to convert the $f$ values to $\kappa$, the spectra of all the three scalars (temperature, moisture, and $\rm CO_{2}$) and streamwise velocity ($u$) follow a $-1$ spectral slope at lower wavenumbers (Fig. 7a). Regarding $w$, the $-1$ slope is not clearly evident, which is consistent with the attached eddy hypothesis (Townsend, 1976). For the scalars, especially moisture and $\rm CO_{2}$, one could argue that at larger wavenumbers the spectra do not exactly follow the $-5/3$ slope, which could be due to attenuation of the high frequency fluctuations in the gas analyzer sampling tube and instrument noise. However, under the transformation with Eq. 11 ($\kappa z=2\pi fz/U_{C}(\kappa z)$), an extensive region of $-5/3$ slope (nearly two decades, $10.01<\kappa z<$ 1) is recovered for $u$ and the scalars (Fig. 7b).
In fact, Eq. 11 modifies the wavenumber ranges up to which the scaling laws hold. For instance, it shrinks the extent of $-1$ spectral slope while inflating the regions of $-5/3$ (see Fig. 7b). Although in Fig. 7b we show the ensemble-averaged results, if one looks at individual 30-min periods then too this extended $-5/3$ region is conspicuous (see Fig. S5 in the Supplementary Material). For illustration purposes, $u$ and $T$ signals are used as examples in Fig. S5. Upon applying the Kolmogorov’s similarity model on these transformed spectra, one could essentially recompute the dissipation rates of turbulence kinetic energy and scalar variances. This is in accord with previous studies, demonstrating that the dissipation rate estimations from temporal turbulence spectra are sensitive to whether the Taylor’s hypothesis holds (Dahm and Southerland, 1997; Hsieh and Katul, 1997).
Unlike scalars, for the vertical velocity, scale-dependent convective speeds modify the $-5/3$ spectral slope. We explain this anomaly in Appendix C. If we turn our attention towards the flux cospectra (heat, moisture, momentum, and $\rm CO_{2}$), then under the assumption of Taylor’s hypothesis, the cospectral amplitudes at larger wavenumbers show a slope that is nearly equal to $-7/3$ (Fig. 7c). The cospectral slope of $-7/3$ is consistent with the dimensional result of Wyngaard and Coté (1972). For momentum and $\rm CO_{2}$ transport we show their absolute cospectral densities, since their original values are negative. This is evident from the premultiplied cospectra shown in Fig. S6 of the Supplementary Material. Now, if the conversion to wavenumbers is carried out using Eq. 11, we note that the flux cospectral densities display a $-5/3$ scaling instead of $-7/3$ (Fig. 7d). The exact cospectral slope has been debatable since their exponents have shown to be not unique and depend on how the dimensional analysis is performed (Cheng et al., 2020). These same results appear to be true even for the perpendicular-wind cases (Figs. 7e–h). Inspite of such interesting findings, the estimations of the revised $\kappa z$ values and subsequently the spectra are hinged upon how accurate the DTS measurements are. While we discuss this aspect next, we also provide a recipe on how to correct the EC flux measurements based on the results presented so far.
3.3 Implications for EC fluxes
3.3.1 Limitations of DTS
Peltola et al. (2021) showed that at smaller scales of the flow the DTS measurements could be affected by the presence of a white noise. Since white noise has a flat spectrum, in a premultiplied form, it displays a $+1$ spectral slope. To check this, we first computed the DTS spectra of temperature fluctuations from temporal data, and compared the same with EC measurements. To aid the comparison, the spectral amplitudes are normalized by the temperature variances ($\sigma^{2}_{T}$) estimated from the EC measurements and the frequencies are scaled with the standard surface layer scaling ($fz/U$) of Kaimal et al. (1972). Here, we focus only on the along-wind periods although the conclusions remain same for the perpendicular-wind periods too. From Fig. 8a, one could see that the DTS temporal spectra agree exceptionally well with the EC temperature spectra at lower frequencies. However, since the temporal resolution of DTS data ($2.1\ \rm s$) is lower than the EC one ($0.1\ \rm s$), it does not fully capture the fine scale features. Contrary to Peltola et al. (2021), we do not find any evidence of a white noise since there is no presence of a $+1$ spectral slope at larger frequencies. This is likely since our analyses focused on periods with high temperature variance and hence on periods with high signal-to-noise ratio.
Figure 8: For the along-wind cases, (a) the ensemble-averaged pre-multiplied temporal spectra of temperature ($fS_{TT}(f)$) from the DTS data are compared against the EC measurements of temperature. The spectral amplitudes are normalized with the variance of EC temperature ($\sigma^{2}_{T}$) while the frequencies are scaled with the canonical surface layer scaling. (b) The spatial spectra from the DTS data are shown for the along-wind scenario. The $-2/3$ and $+1$ slopes denote the predictions from the Kolmogorov’s spectral model and a pure white-noise case. The vertical dash-dotted line indicates the location $\kappa z=3$. (c)–(d) The impact of using scale-dependent convection speeds instead of a constant one is shown for the heat and moisture cospectra. Here, the scale-dependent convection speeds are estimated from our DTS measurements (pink lines) and from the Cheng et al. (2017) data (blue lines). The red and magenta lines denote when the conversion was carried out using mean wind speed from the EC data ($U$) and bulk convective speed ($U_{C}$) estimated from the space-time correlation curves. (e) For all such different convection speeds, frequencies are plotted against the converted wavenumbers. The horizontal dash-dotted line indicates the critical wavenumber $2\pi/0.3\rm\ m^{-1}$ while the vertical one indicates $f_{\rm cr}$ (Eq. 15). (f) The scatter plot between the surface energy balance components is shown. Here, $\rm R_{n}$ and $\rm G_{0}$ are the net-radiation and soil heat flux components while $\rm SHF$ and $\rm LHF$ represent the sensible and latent heat fluxes in $\rm W\ m^{-2}$. The red squares represent the original values of the sensible and latent heat fluxes and the blue dash-dotted line indicates the best fit. The pink squares represent the values after applying Eq. 17 on the sensible and latent heat fluxes.
On the other hand, if the same information is seen from spatial domain (Fig. 8b), then the DTS premultiplied temperature spectra do show a clear $-2/3$ spectral slope up to wavenumbers $\kappa z\leq 3$ (the vertical dash-dotted line in Fig. 8b). Beyond that, a spurious increase is noted in the spectra although that increase cannot be entirely described as a pure white noise. Nevertheless, this raises a question about the accuracy of $U_{C}(\kappa z)$ values in Fig. 5a. To check that, we carried out a randomly-shuffled exercise on the DTS temperature data. At each instant of time, we randomly shuffled the spatial DTS data, thereby destroying any coherence between them. For randomly-shuffled DTS data, the phase differences between the spatial temperature signals at time instants $t$ and $t+\Delta t$ would be equal to $\pi/2$ at all the wavenumbers since no coherence is expected between them. Therefore, Eq. 5 suggests that for randomly-shuffled signals, $U_{C}(\kappa)$ would vary as ${(\kappa)}^{-1}$, since both $\varphi$ and $\Delta t$ remain constant irrespective of different $\kappa$ values.
This is indeed the case as shown in Fig. S7 of the Supplementary Material. As opposed to the real signals (Fig. S7a), for the randomly-shuffled samples, one can clearly see that the $U_{c}(\kappa z)$ values vary as ${(\kappa z)}^{-1}$ instead of ${(\kappa z)}^{-0.8}$ (Fig. S7b). For clarity purposes, we use a linear $x$ axis in Fig. S7 and restrict the wavenumbers up to $\kappa z\leq 3$. This outcome provides confidence in our estimates of $U_{c}(\kappa z)$ in Fig. 5a. We also note that the exponent $-0.8$ is relatively close to $-1$ and hence our results can be partly affected by the noise seen in Fig. 8b. Nevertheless, in Fig. S8 of the Supplementary Material, we compare three scaling relationships (${(\kappa z)}^{-1/3}$, ${(\kappa z)}^{-0.8}$, and ${(\kappa z)}^{-1}$) for the perpendicular-wind cases, and it is conspicuous that at wavenumbers $\kappa z>1$, $U_{C}(\kappa z)$ values are best described by the relationship ${(\kappa z)}^{-0.8}$.
3.3.2 A critical frequency limit
Despite the limitations of DTS measurements, we test one hypothesis of Cheng et al. (2017) where they proposed that by using scale-dependent convection speeds one can increase the flux contributions from smaller time scales. This eventually could serve as a correction for improving the imbalance in surface energy components as commonly being observed in micrometeorological measurements (Mauder et al., 2020). In contrast to ours, Cheng et al. (2017) found that the scale-dependent convection speeds of temperature structures vary as ${(\kappa z)}^{-1/3}$. If we use that information to convert the frequencies to wavenumbers, then after some algebraic manipulation, one can show that the wavenumbers ($\kappa z$) can be estimated from frequencies ($f$) by solving the polynomial,
${(\kappa z)}^{3}=a(1+\kappa z),$ (13)
where $a=(2\pi fz/U_{C})^{3}$. To derive Eq. 13, we have assumed that the convection speeds can be expressed as,
$\frac{U_{c}(\kappa z)}{U_{c}}=\frac{1}{{(1+\kappa z)}^{1/3}},$ (14)
which appears to be supported by the plot shown in Fig. S4b of the Supplementary Material.
In Figs. 8c-d, we show the premultiplied heat and moisture cospectra ($fS_{wx}(f)$, where $x=T,q$) where the frequencies are converted to wavenumbers using mean wind speed ($U$, red line), bulk convective speed ($U_{C}$, magenta line), and scale-dependent convective speeds ($U_{C}(\kappa z)$) from our observations (pink line) and from Cheng et al. (2017) (blue line). We adopt the premultiplied form since the area under this log-linear coordinate system is directly related to the flux contributions across different scales (Kaimal and Finnigan, 1994). The results from Figs. 8c-d clearly demonstrate that using our scale-dependent convective speeds (Eq. 9) cause an appreciable increase in the heat and moisture fluxes at smaller scales of the flow. More importantly, the differences between the cospectral shapes when using either $U$, $U_{C}$, or $U_{C}(\kappa z)$ relationship from Cheng et al. (2017) to convert the frequencies to wavenumbers are small.
The transformations between wavenumbers and frequencies (Eqs. 11 and 13) also yield an unrealistic outcome since the wavenumbers estimated from both Eqs. 11 and 13 are strong non-linear functions of frequency. If these functions are plotted for a specific 30-min period, as done in Fig. 8e, it is apparent that the wavenumbers estimated from Eq. 11 can be up to $10^{6}\rm\ m^{-1}$ when converting the EC sampling frequency (10 Hz) to wavenumber. This cannot be true as it suggests that the EC systems can measure eddies comparable to the dissipation scales, which are clearly smaller than the path length (distance between the transducers) of the sonic anemometers themselves. For our sonic anemometer system, the path length is around 16 cm, and hence the wavenumbers that can be realistically resolved is around $2\pi/0.3\rm\ m^{-1}$. To avoid the aliasing effect, the minimum wavelength being resolved is assumed to be twice the path length (i.e. 30 cm). This limit is shown as a horizontal dash-dotted line in Fig. 8e. If we call this minimum wavelength being set by the physical dimension of the sonic anemometer as $\kappa_{\rm cr}$, then from Eq. 11, one can convert $\kappa_{\rm cr}$ to a frequency limit ($f_{\rm cr}$) in the temporal space. Accordingly, one can estimate $f_{\rm cr}$ as,
$f_{\rm cr}=\frac{\kappa_{\rm cr}U_{C}}{2\pi{(1+\kappa_{\rm cr}z)}^{4/5}}.$ (15)
For our set up, $z=3.1$ m, $\kappa_{\rm cr}=2\pi/0.3\rm\ m^{-1}$, and $U_{C}$ can be found from the space-time correlation curves (see Fig. 4). For the along-wind cases, $U_{C}$ values are averaged over all the five sections. By replacing all these values in Eq. 15, $f_{\rm cr}$ appears to be $\approx$ 0.3 Hz (for both along- and perpendicular-wind periods), with some run-to-run variations since $U_{C}$ values are not fixed.
3.3.3 High-frequency flux uncertainties
If we consider the $U_{C}(\kappa z)$ values and therefore the derived wavenumbers from Eq. 11 to be correct, then it is unclear what the EC observations represent at frequencies above $f_{\rm cr}$ (i.e. $f\geq f_{\rm cr}$). This is because, beyond $f_{\rm cr}$, the eddy length scales are smaller than the physical dimension of the instrument itself. Of course this problem is not that severe if the wavenumbers are estimated from Eq. 13, i.e. using the Cheng et al. (2017) relationship (see the blue line in Fig. 8e). By following the same procedure, $f_{\rm cr}$ values for the Cheng et al. (2017) relationship can be derived as,
$f_{\rm cr}=\frac{\kappa_{\rm cr}U_{C}}{2\pi{(1+\kappa_{\rm cr}z)}^{1/3}}.$ (16)
If one substitutes all the values from our set up in Eq. 16, $f_{\rm cr}$ appears to be $\approx$ 2 Hz. In general, $f_{\rm cr}$ values are dependent on the power-law exponent ($\alpha$) of the scale-dependent convection speeds, $U_{C}(\kappa z)/U_{C}=1/(1+\kappa z)^{\alpha}$, with $10<\alpha<$ 1. Here, $\alpha=1$ corresponds to a limit for the randomly-shuffled samples while $\alpha=0$ indicates no scale-dependency.
The existence of $f_{\rm cr}$ ultimately results in a situation where the fluxes associated with EC systems at higher frequencies are largely uncertain. In a fractional form, this high-frequency flux uncertainty can be expressed as,
[…]
where $S_{wx}$ is the cospectral density between $w$ and $x$. Here $x$ can be any constituent, e.g. $u$, $T$, H2O or CO2. For any EC system, this fraction can be estimated by computing $f_{\rm cr}$ from either Eq. 15 or Eq. 16, depending on the flow situations. In the absence of any concrete evidence, we hypothesize that Eq. 15 is valid for highly-heterogeneous flows (e.g., forest clearcut) while Eq. 16 can be used for homogeneous surface layer flows. For highly-heterogeneous conditions, $U_{C}$ can be obtained from Eq. 7.
For our case, the fraction of fluxes within the frequency range $0<f<f_{\rm cr}$ is around 85-90% of the total flux being resolved. Thus, between 10-15% of the flux is related to the frequency range that is questionable given the physical dimension of the sonic-anemometer (see above). This has significant impact on the long-term estimations of ecosystem gross primary productivity, net ecosystem exchange, evapotranspiration, sensible heat and momentum fluxes from the EC systems. Potentially, Eq. 17 could be extended to reevaluate the estimations of velocity and scalar variances and turbulence kinetic energy, and therefore has implications for surface layer parameterizations. One other impact of Eq. 17 is on the surface energy balance.
[…]
From a practical standpoint, since all our results are from a single height, we cannot establish whether $z$ is the ideal scaling variable to describe $U_{c}(\kappa)$. If $z$ indeed is the correct scaling parameter, then depending on the observation height, $f_{\rm cr}$ values will change (see Eqs. 15 and 16). Accordingly, the high-frequency flux uncertainty associated with the EC systems would also change, and possibly will increase at higher heights since $f_{\rm cr}$ is inversely proportional to $z$. We present our conclusions in the next section.
4 Conclusion
In this study, we have developed a comprehensive framework to test the validity of Taylor’s hypothesis and its impact on turbulence statistics over a highly heterogeneous forest clearcut surface. To execute this objective, an extensive dataset encompassing both Distributed Temperature Sensing (DTS) and Eddy Covariance (EC) measurements were used. These datasets were collected during a five-month field campaign (May-September 2024) and were separated into two different categories. First, when the mean wind speed was parallel to the forest edge, and second, when it was perpendicular to the edge. For the former, our results demonstrated that the Taylor’s hypothesis remained invalid as the turbulent intensities of the streamwise wind fluctuations exceeded 0.5. An empirical relationship is proposed to estimate the convective speeds in such extreme conditions from the sonic anemometer observations. Corresponding to the latter scenario, no such relationship was apparent since depending on whether the wind approached from the forest or from the clearcut the results were different.
Despite this discrepancy, for both along- and perpendicular-wind cases, the ratios of the convective speeds of positive and negative temperature structures remained at a constant value of 0.8, irrespective of the temperature skewness measured by the EC system. We used this information to compute the sizes of up- and down-plumes in a convective boundary layer. We found that the sizes of the down-plumes can be as large as 12 times the measurement height and therefore would require considerably larger area for their statistics to converge. This result informs about what should be the appropriate horizontal domain size while setting up the large eddy simulations of convective boundary layers. Furthermore, the different convection speeds for the positive and negative temperature structures suggests that an EC tower sampling the flow over time produces a biased sample of the flow when compared to sampling taken over space.
We demonstrated that the convective speeds of temperature structures at inertial subrange scales behave in a power-law fashion with increasing wavenumbers. These wavenumbers were directly computed from the DTS spatial data. Contrary to a previous study over a homogeneous atmospheric surface layer flow (Cheng et al., 2017), the exponent of this power law was $-0.8$ instead of $-1/3$. The $-1/3$ exponent was theoretically derived under the assumption of homogeneous and isotropic turbulence at smaller scales of the flow. For the present case in hand, we hypothesize that the heterogeneity in the surface conditions violate these assumptions even at smaller scales of the flow. We evaluated the impact of this revised scale-dependent convection speed on the turbulence spectra of velocity and scalars (temperature, moisture, and $\rm CO_{2}$), estimated from the temporal EC measurements. The results displayed an extensive $-5/3$ spectral slope (over two decades of wavenumbers) for the streamwise velocity component and the three different scalars when the frequencies were converted to wavenumbers using our scale-dependent convective speeds. For the flux cospectra, instead of a $-7/3$ slope, a $-5/3$ slope was recovered under such conversion.
The appearance of a slope flatter than $-7/3$ indicates an enhancement of turbulent fluxes at smaller scales of the flow. However, the non-linear conversion of frequencies to wavenumbers through scale-dependent convective speeds generates unrealistically large wavenumber values. The physical dimension of the sonic anemometers sets a strict constraint on the wavenumbers that can be resolved by the EC systems. In a temporal domain, this eventually results in a critical frequency limit beyond which the EC measurements cannot be trusted. This critical limit depends on the measurement height and on the power-law exponent of the scale-dependent convective speeds. The existence of this limit dictates that there is a high-frequency flux uncertainty associated with EC systems, which, in turn, has an adverse effect on the estimation of surface energy budget. Before we outline our future research directions, Hilland and Christen (2024) asked whether the Taylor’s hypothesis is ever satisfied in atmospheric flows. For a flat desert surface, Hilland and Christen (2024) found that the Taylor’s hypothesis appeared to be valid only between spatial scales 5 to 16 m. Over a forest clearcut, our results suggest that for the periods when the wind approached the clearcut location from the forest, the Taylor’s hypothesis remained valid at scales $\kappa z<1$, where $\kappa$ is the streamwise wavenumber and $z$ is the measurement height. This was realized notwithstanding the fact that during such periods the flow was disturbed as it encountered abrupt changes in the surface roughness.
At present, these results are representative of buoyancy-driven flows in a heterogeneous location and are also limited by a single observation height. One another drawback is, the estimations of the convective speeds are based on an imperfect temperature signal and we inherently assume that the same could be extended to other scalars and velocity components. It is not immediately apparent whether the structures in all these different turbulent variables move at same speeds. Therefore, for future studies, more extensive spatial data are needed to verify our framework, which ideally should span multiple heights, stability conditions (stable, neutral, and convective), surface roughness (forest canopies, grasslands, arid deserts, water-bodies, etc.), and signal types (velocity and scalars). To achieve this ambitious goal, the observations alone will not suffice and large eddy simulations would be needed to complement the DTS measurements.
Acknowledgements.
[…]
Appendix A Assessment of stability
Figure 9: For the along-wind half hour periods from the EC clearcut measurements, (a) the PDFs of velocity and temperature fluctuations are shown. The green dash-dotted line indicates the Gaussian distribution, while the pink dash-dotted line represents the Gram-Charlier distribution. (b) The JPDFs between the vertical velocity and temperature fluctuations are shown where the colorbar represents the contour levels. (c) The flux-weighted JPDFs between the same two signals are shown. (d) The scatter plot between the correlation coefficients $R_{wT}$ and $R_{wu}$ are shown for both along and perpendicular wind cases with respect to the forest edge. The absolute values of $R_{wu}$ ($|R_{wu}|$) are used and the dash-dotted line indicates the 1:1 relationship.
In this appendix, we report the statistics of velocity and temperature fluctuations from the EC measurements corresponding to the along-wind conditions. The conclusions remain the same for perpendicular-wind conditions and hence not being shown. These statistics are used to assess the stability conditions. In practice, the Obukhov similarity parameter is commonly used to determine the atmospheric stability, but we deliberately refrained from that approach since this theory was developed for homogeneous flows and its validity is questionable over heterogeneous surfaces such as a forest clearcut. In Fig. 9a, we show the PDFs of velocity and temperature fluctuations. Typical to a buoyancy-driven flow, the PDFs of temperature fluctuations remain strongly non-Gaussian as opposed to the velocity ones ($u^{\prime}$ and $w^{\prime}$). In fact, the PDFs of temperature fluctuations can be represented through a cumulant expansion method, otherwise known as Gram-Charlier distribution (Chowdhuri and Prabha, 2019). This distribution can be expressed as:
$P(T^{\prime}/\sigma_{T})=\frac{1}{\sqrt{2\pi}}\exp[-{(\frac{T^{\prime}}{\sigma_{T}})}^{2}][1+\frac{\mathcal{S}(T)}{6}(\mathcal{S}(T)-3\frac{T^{\prime}}{\sigma_{T}})],$ (18)
where $\mathcal{S}(T)$ is the skewness of temperature fluctuations. The non-Gaussianity in temperature statistics is also reflected in how the JPDFs between the vertical velocity and temperature fluctuations behave. From Fig. 9b, it is apparent that the JPDFs are strongly skewed towards quadrant III, which represents the downdraft motions. To further investigate the heat transport characteristics for our selected runs, we computed the flux-weighted JPDFs, expressed as, $w^{\prime}T^{\prime}P(w^{\prime},T^{\prime})$, where $P(w^{\prime},T^{\prime})$ is the JPDF between $w^{\prime}$ and $T^{\prime}$. The contours of this quantity are plotted in Fig. 9c, and the area under them is proportional to the amount of heat being transported (Nakagawa and Nezu, 1977). It is immediately clear from Fig. 9c that the transport of heat is strongly dominated by the down-gradient quadrants (I and III) with very small amounts being transported in counter-gradient directions (II and IV).
This organization results in a situation where the heat is transported more efficiently than momentum. One could verify this argument in Fig. 9d, where the correlation coefficients between $w^{\prime}$ and $T^{\prime}$ ($R_{wT}$) remain considerably larger than the ones between $w^{\prime}$ and $u^{\prime}$ ($R_{wu}$). Since $R_{wu}$ values are negative, their absolutes ($|R_{wu}|$) are used on the $x$ axis of Fig. 9d. Additionally, the $R_{wT}$ values are nearly constant at around $0.48$, which is very close to $0.52$ and can be derived from free-convective scaling relations for the vertical velocity and temperature fluctuations (Wyngaard, 2010). Therefore, by combining all these insights, one can claim that our selected periods were prototypical of a buoyancy-driven flow.
Appendix B Applicability of elliptical model
Figure 10: For a specific 30-min period (10:30-11:00 UTC, 18 May 2024) and Section-1 DTS measurements, (a) the contours of space-time correlation curves are shown where the $x$ and $y$ axes represent the temporal and spatial separations, respectively. The colorbar denotes the correlation values. (b) The space-time correlation curves are shown for two specific $r$ values in order to illustrate the computation of $\tau_{0}$. (c) The scatter plots between the separation distance $r$ and $\tau_{0}$ and $\tau_{p}$ are shown, where the straight line fits are used to compute the sweeping and convection velocities separately. (d)–(e) The scatter plots between the mean wind speed from the EC measurements and the convection and sweeping velocities are shown. The different gray colored points indicate the values from the five different horizontal sections of the DTS set-up. The dash-dotted black line indicates the 1:1 relationship. (f) The scatter plots between the turbulence intensities from EC measurements and ratios of the resultant convection speed with EC mean wind are shown for the five different horizontal sections.
As opposed to a linear model to estimate the convective speeds of turbulent structures, there also exists an elliptical model. A couple of previous studies, pertaining to atmospheric surface layer and sparse canopy flows, found the elliptical model to be more appropriate while estimating the convective speeds of turbulent structures (Han and Zhang, 2022; Everard et al., 2021). In this model, one assumes the contours of space-time correlation curves can be approximated as an ellipsoid. To test this model for a highly-heterogeneous flow, such as here, in Fig. 10a we show the contours of space-time correlation curves for the period between 10:30-11:00 UTC, corresponding to Section-1 of the DTS cable. From a visual inspection, one could say that the contours of $R_{TT}(r,\tau)$ appear to be of elliptical shape. To quantify this aspect, the elliptical model computes two velocities $U_{e}$ and $V_{e}$, which are referred to as convection and sweeping velocities, respectively.
For all practical purposes, to estimate these two quantities one needs to use two different timescales (Han and Zhang, 2022). One of those timescales is $\tau_{0}$ and are calculated from the condition, $R_{TT}(r,0)=R_{TT}(0,\tau_{0})$. The second one is $\tau_{p}$, which represents the peak positions of the space-time correlation curves at different values of $r$. The concepts of $\tau_{0}$ and $\tau_{p}$ are illustrated in Fig. 10b. The blue square represents the peak of the space-time correlation curve at a specific value of $r$, thereby denoting $\tau_{p}$. On the other hand, $\tau_{0}$ represents that timescale where $R_{TT}(r,0)$ is equal to $R_{TT}(0,\tau_{0})$. For this specific example, $r=21\ \rm m$. After defining those, one fits straight lines to quantities $\tau_{0}$ and $r$, and $\tau_{p}$ and $r$ (see Fig. 10c). If the slopes of these lines are $\alpha_{0}$ ($\tau_{0}=\alpha_{0}r$) and $\beta_{p}$ ($\tau_{p}=\beta_{p}r$), respectively, then $U_{e}$ and $V_{e}$ are computed as,
$U_{e}=\frac{\beta_{p}}{\alpha^{2}_{0}},V_{e}=\frac{\sqrt{1-\beta^{2}_{p}/\alpha^{2}_{0}}}{\alpha_{0}}.$ (19)
In Figs. 10d-e, we compare both $U_{e}$ and $V_{e}$ with the mean wind speed ($U$) from the EC measurements. This comparison is done for the along-wind cases where the gray-colored squares represent the five different sections of the DTS cable. It is clear from Fig. 10d that $U_{e}$ is weakly correlated to $U$, which is in stark contrast with Fig. 4a. The sweeping velocities, on the other hand, are consistently smaller than $U$ (Fig. 10e). In an elliptical model, the resultant convective speeds are defined as, $\sqrt{U_{e}^{2}+V_{e}^{2}}$ and if this ratio $\sqrt{U_{e}^{2}+V_{e}^{2}}/U$ is compared against the turbulent intensities of the streamwise velocity component ($\sigma_{u}/U$), a significant scatter is observed (Fig. 10f). In fact, the relationship as described by Eq. 7, cannot fully explain these variations. We, therefore, conclude that for our situation the performance of the elliptical model remains poor. It is, however, unclear whether this is a general feature of heterogeneous flows or a feature specific to our site. More experimental validation is needed to confirm this aspect, which is beyond the scope of the present work.
Appendix C Spectral conversion
Figure 11: The parametric forms of vertical velocity spectra are shown with respect to (a) frequencies (Eq. 20) and (b) wavenumbers (Eq. 22). (c)–(d) The ensemble-averaged vertical velocity spectra of Figs. 7a–b are replotted. The gray and black dash-dotted lines denote the spectral slopes $-1.25$ and $-5/3$, respectively.
We derive an analytical expression to convert the spectrum from temporal to spatial domain by using scale-dependent convective speeds. We consider the spectrum of vertical velocity, which can be expressed through the following parametric form (Kaimal and Finnigan, 1994),
$S_{ww}(f)=\frac{a}{{(1+bf)}^{5/3}}.$ (20)
[…]
$f$ $=\frac{U_{c}\kappa}{2\pi{(1+\kappa z)}^{4/5}}$ (21)
$df$ $=\frac{U_{c}(5+\kappa z)}{10\pi{(1+\kappa z)}^{9/5}}d\kappa,$
where $U_{C}$ is the bulk convective speed and $z$ is the measurement height (3.1 m). By substituting them in Eq. 20, we obtain the following spectrum in spatial domain,
$S_{ww}(\kappa)=\frac{aU_{C}{(2\pi)}^{5/3}(5+\kappa z)}{10\pi{(1+\kappa z)}^{7/15}{(2\pi{(1+\kappa z)}^{4/5}+b\kappa U_{C})}^{5/3}}.$ (22)
For the same 30-min run, we show this parametric spectrum in Fig. 11b. In contrast to Fig. 11a, Eq. 22 modifies the $-5/3$ spectral slope with a one equal to $-1.25$. Therefore, the scale-dependent convective speeds do not yield the desired result for the vertical velocity.
This conclusion is even supported by the real data. For illustration purposes, corresponding to the along-wind cases, we plot the ensemble-averaged vertical velocity spectra in Figs. 11c–d. In Fig. 11c, the frequencies are converted to wavenumbers through a linear transformation, i.e. using a constant wind speed ($U$). On the other hand, in Fig. 11d, the conversion is carried out with scale-dependent convective speeds ($U_{C}(\kappa z)$). Since the wavenumbers and spectral densities are normalized in Figs. 11c–d, their ranges are different than Figs. 11a–b. Similar to Fig. 11b, in Fig. 11d, the conversion with scale-dependent convective speeds impacts the $-5/3$ spectral slope and replaces it with $-1.25$. Instead, one recovers a $-5/3$ scaling with a constant convective speed (Fig. 11c). Contrary to $w$, the scalars show a clear $-5/3$ spectral slope when the scale-dependent convective speeds are used (Fig. 7b). Thus, in terms of how fast the eddies move, dissimilarities exist between the scalars and vertical velocities. At present, this is a hypothesis and needs to be confirmed through large eddy simulations.