### v2

# Taylor’s hypothesis and its impact on turbulence measurements in a forest clearcut

Abstract
Taylor’s hypothesis converts observations done over time to spatial information of the flow while carrying out measurements on a micrometeorological tower. To address its validity over a highly heterogeneous forest clearcut surface, we utilize an extensive Distributed Temperature Sensing (DTS) and Eddy Covariance (EC) datasets. The DTS measured space-time correlation curves of temperature fluctuations are used to compute the bulk convective speeds of temperature structures ($U_{C}$) in buoyant and highly turbulent conditions at a height of $\approx$ 3.1 m above the clearing. These convective speeds are compared with the mean wind speed ($U$) and depending on if $U$ is parallel or perpendicular to the forest edge, the relationships show differences. Particularly, for the parallel-wind scenario, $U_{C}>U$, which we explain by invoking sweeping effects. However, irrespective of the wind direction, the convective speeds of temperature structures at inertial subrange scales behave in a power-law fashion (having an exponent of 0.8-0.8) with increasing wavenumbers. These scale-dependent convective speeds impact the properties of temporal turbulence spectrum in the inertial subrange scales by yielding slopes larger than what one would obtain through Taylor’s hypothesis. In our highly turbulent case, these slopes appear enhanced than 5/3-5/3, and we provide a plausible explanation for this by conducting a dimensional analysis. Moreover, the scale-dependent convective speeds non-linearly transform the temporal frequencies to streamwise wavenumbers. This yields a critical frequency limit, beyond which the eddy length scales derived from frequencies are smaller than the physical dimension of the sonic anemometers, and therefore cannot be faithfully resolved.
[…]
The measurements of near-surface turbulent flows serve as a cornerstone in land-atmosphere interaction research. Sans spatial information, the accuracy of these turbulence measurements is tied to Taylor’s hypothesis, which states that the space-time conversion can be accomplished through the local mean wind speed. We test Taylor’s hypothesis over a highly heterogeneous forest clearcut surface by using fiber-optic cables that acquired air temperatures at high spatial and temporal resolutions. From this data, we estimate the speeds of turbulent eddies at each spatial scale of the flow and show how that modifies the turbulence statistics when computed over temporal domain. Due to a non-linear transformation of time to space, a critical frequency limit exists beyond which the space-time conversion appears uncertain.
1 Introduction
The quantification of vertical turbulent transport of heat, matter (e.g. moisture, $\rm CO_{2}$) and momentum are important for understanding the land-atmosphere interaction, energy budget of the Earth’s surface, and carbon and water exchanges between the ecosystem and atmosphere. Under the assumptions of homogeneity and stationary conditions, these vertical transports can be interpreted as surface fluxes, which eventually serve as boundary conditions for the Earth System Models. Given such various applications, the micrometeorological towers at different locations across the globe routinely monitor these fluxes under the aegis of various programs such as FLUXNET, AmeriFlux, NEON, and ICOS (baldocchi2001fluxnet; novick2018ameriflux; metzger2019neon; heiskanen_integrated_2022). Although the turbulent eddies facilitating the vertical transport exist physically in space, spatial information of atmospheric turbulent flows are rarely available and thus the instruments mounted on micrometeorological towers register the turbulent variables over time (hilland2024systematic). To circumvent this issue, Taylor’s frozen turbulence hypothesis is used to convert the temporal information to spatial ones. This hypothesis states that the turbulent structures of any spatial dimension move past the measurement location with a convection velocity in a frozen state, i.e. their characteristics change slowly compared to the speed at which the mean flow transports them along. In the absence of spatial information, it remains experimentally challenging to verify the accuracy of the Taylor’s hypothesis (TH) under various land-surface conditions.
Despite so, a handful of previous studies attempted to verify TH either through theoretical formulations or by conducting novel field experiments that provided spatial information. From the theoretical side, the early study of wyngaard1977taylor used a model developed by lumley1965interpretation to investigate the effects of TH on the turbulence statistics at smaller scales of the flow. In particular, wyngaard1977taylor provided theoretical formulations that accounted for the deviation from the assumption of constant convective speeds across all scales of the flow. Later on, tong1996taylor experimentally verified their results by using two spatially separated ultrasonic anemometers. Except a few experimental set-ups, spatially explicit sonic anemometer measurements are not readily available (han2019applicability). Some studies deployed thermocouples to test TH but their spatial resolution was rather limited (deardorff1982investigation; everard2021sweeping). However, with the advent of novel instrumentation, spatial data of turbulent variables can be obtained at high enough resolution either from remote sensing or through fiber optic cables. In that regard, higgins2012effect used a Raman Lidar to investigate the convective speeds of temperature and moisture structures over a landscape surrounding a water body. On the other hand, cheng2017failure used distributed temperature sensing (DTS) measurements over a grassland to verify the hypothesis for different atmospheric stability conditions.
The DTS measurement technique uses fiber optic cables and relies on Raman scattering principle to convert the back-scattered energy from the emitted light pulses to temperature values at specific locations in space (peltola2021suitability). By laying out these cables either horizontally or vertically, it is therefore possible to obtain temperature data at a fine enough spatial (of the order of tens of cm) and temporal resolutions (of the order of seconds). cheng2017failure used this spatial temperature data from a horizontal configuration to evaluate scale-dependent convective speeds of turbulent structures. For their experimental dataset over a flat grass surface, they found that at smaller scales of the flow, the turbulent structures moved at speeds that scaled with the streamwise wavenumbers ($\kappa$) as ${(\kappa)}^{-1/3}$. To reach this conclusion, cheng2017failure adopted the technique of del2009estimation, where differences in Fourier phases of spatial temperature signals at two subsequent instant of time were used to estimate the wavenumber-dependent convection speeds. cheng2017failure also provided theoretical support to this 1/3-1/3 exponent by borrowing formulations from the literature of homogeneous and isotropic turbulence. They hypothesized that these scale-dependent convective speeds could enhance the turbulent fluxes by 10-20% at smaller scales of the flow, which, eventually, could solve the crisis of surface energy imbalance.
In recent times, DTS measurements were used over a flat and homogeneous Namib desert surface to verify TH (hilland2024systematic). hilland2024systematic found that the TH was almost never satisfied in atmospheric surface layer flows under various stability conditions (ranging from buoyancy- to shear-dominated flows), and the measurement height was an important parameter in that regard. Contrary to cheng2017failure, hilland2024systematic used a correlation-cube methodology to estimate the scale-dependent convective speeds. In their method, the definition of scales is not related to streamwise wavenumbers in a Fourier domain. In spite of this, hilland2024systematic found that the convective speeds of smaller-scale turbulent structures were indeed lower than the bulk convective speeds of larger-scale structures. Be that as it may, hilland2024systematic did not report any power-law variations in their estimates of scale-dependent convective speeds.
All these studies described above were conducted over homogeneous atmospheric surface layer flows. However, in reality, one mostly encounter strong heterogeneity in land-surface conditions, which presents a significant challenge to the atmospheric community (bou2020persistent). One of such examples is a forest clearcut. Clearcutting is a widely utilized forest management practice to collect timber and in some cases also to restrict the spread of wildfires by limiting the access to the fuel (ortega2024modeling). From a fluid dynamical perspective, clearcutting introduces additional complexities to the air flows. For instance, when the air blows directly from the forest canopy sub-layers to clearcut location, it encounters a significant change in surface roughness, which causes the flow to deviate from being in equilibrium with local surface characteristics. The laboratory experiments of chung2021turbulence over an artificial heterogeneous canopy showed that these deviations are often caused by the canopy-scale eddies extending their footprints on the clearcut surface depending on the distance from the canopy edge. On the other hand, the field-experimental results of detto2008structure provided evidence that the flow near the forest edge shares several attributes with a backward-facing step flow in which a recirculation zone dominates the turbulence statistics.
From a climate science standpoint, the forests are strong sources and sinks of water vapor and carbon dioxide, respectively, thereby influencing the water- and carbon-cycles between the atmosphere and the ecosystem. Forest clearcutting clearly impacts these exchange processes and the climatic impact of such forest management practices can be observed with the eddy covariance (EC) technique (tikkasalo2025eddy). However, due to the aforementioned complexities, the turbulent behavior of the air flow in forest clearcuts pose challenges while interpreting the observed flux values. One of such aspects is regarding the applicability of TH. For instance, if the convective speeds of turbulent structures are indeed scale-dependent then it can have serious consequences on the EC flux measurements as hypothesized by cheng2017failure. Furthermore, the relation derived by cheng2017failure to describe the scale-dependent convective speeds was under the assumption of homogeneous and isotropic turbulence. A recent study by chowdhuri2024quantifying revealed that the turbulence is hardly isotropic at smaller scales of the flow corresponding to the heights closer to the surface. Moreover, it is uncertain whether the assumption of homogeneity would remain valid in highly heterogeneous conditions, such as in a forest clearcut.
In many respects, forest clearcut flows share resemblance with roughness sublayer flows (flesch_wind_1999). The study by everard2021sweeping on roughness sublayer flows over a vineyard canopy has suggested that when the turbulence intensities are high, strong sweeping effects associated with large eddies of the flow can cause severe modifications to the convective speeds of smaller-scale eddies. This potentially could cause troubles while interpreting the high-frequency EC measurements of turbulence statistics. To address these pertinent issues, our objectives in this study are primarily threefold:
  1. 1.
    Estimate the bulk convective speeds and sweeping speeds of turbulent structures in roughness sublayer flows from the DTS data and use that information to validate TH.
  2. 2.
    Investigate how the convective speeds vary with increasing wavenumbers.
  3. 3.
    Quantify the impact of scale-dependent convective speeds on turbulence statistics and discuss its implications for EC measurements.
To tackle these objectives, we employ an extensive dataset encompassing both DTS and EC measurements, collected during a five-month field campaign in a forest clearcut. In our analysis, we specifically consider two scenarios depending on whether the direction of the mean wind was perpendicular or parallel to the forest edge. In the latter scenario, when the air blows parallel to the forest edge, it does not encounter sharp changes in surface roughness, and therefore, is expected to behave differently than a forest-edge flow. To estimate the convective speeds of the turbulent structures, the space-time correlation curves are constructed from the DTS temperature measurements. This method has been adopted previously for laboratory (deardorff1982investigation), large eddy simulation (su2000two), and field-experimental (everard2021sweeping; han2022applicability) datasets. To compute the sweeping speeds, we follow everard2021sweeping. They showed that the sweeping effects can be quantified by employing the elliptic approximation model of he2010small to describe the space-time correlation curves. In fact, everard2021sweeping provided a recipe to compute the sweeping speeds from field-experimental data. Apart from these computations, we develop a practical framework that accounts for deviations from TH on turbulence spectra and cospectra. This framework could be applied to situations other than a forest clearcut. The rest of this article is organized in three different sections. In Sect. 2 we describe our measurement set up and methodologies being adopted to verify TH. In Sect. 3, we present and discuss our results, and in Sect. 4 future research directions are outlined.
2 Data and Methodology
2.1 Site description
The measurement campaign took place between May 17 and September 10 2024 at Ränskälänkorpi study site (6111 N, 2516 E, 144 m above sea level), which is located on a peatland that was drained before 1960’s for forestry purposes (laurila_set-up_2021). A Norway spruce (Picea abies) stand growing on the site was clear-felled during late winter - spring of 2021 creating a 6.4 ha clearcut area (see Fig. 1) and Norway spruce seedlings were planted after the clearcutting. Few isolated groupings of living and dead trees were left untouched on the clearcut area and logging debris (e.g. branch piles) were left on site. In general, the terrain of the clearcut is flat, but its surface exhibits prominent small-scale undulations due to e.g. drainage ditches and banks. After clearcutting, pioneer species, such as fireweed (Chamaenerion angustifolium), downy birch (Betula pubescens) and raspberry (Rubus idaeus), have spread on the area and created a varying vegetation cover on the clearing (see Fig. 1a). During the study period, the height of these plants varied approximately between 0.5 m and 1.5 m, depending on location and timing of the growing season. In summary, the clearcut area is a complex mosaic of isolated trees, pioneer species, and logging debris creating a rough surface for the flow.
The clearcut is surrounded by a forest dominated by Norway spruce (Picea abies) with height between 20 m and 25 m depending on the location (see Fig. 1b). Leaf area index of the forest in SW direction of the clearing is approximately 2.4 m2 m-2. The forest edges create abrupt changes in canopy height and hence in surface roughness.
Figure 1: (a) Image and a (b) map of the study site. The image was taken towards South-West on June 28 2024 at the location of the white triangle in (b). Different DTS cable sections are highlighted with red lines. In (b) the black contours show the topography with 2-m increments based on the National Land Survey of Finland Topographic Database (data retrieved: 10/2023). The canopy height map was derived from a drone flight conducted during September 27 2022 and complemented with canopy height model of Finnish Forest Center (data retrieved: 12/2024). Note that the canopy height shown in the map does not fully reflect the conditions at the clearing during the study period due to the rapidly developing vegetation cover.
2.2 Measurement setup
The spatial and temporal variability of air temperature above the clearcut was measured with the distributed temperature sensing (DTS) technique thomas_optical_2021. The DTS instrument (Ultima-S, 5 km variant, Silixa Ltd., Hertfordshire, UK) was housed in a temperature controlled cabin in the middle of the forest clearcut (see Fig. 1b). The measurements were conducted using double-ended configuration and signals from both ends of the fiber-optic cable (forward and reverse signals) were stored for later processing (see Sect. 22.3). The DTS instrument provided observations with 1.2 s temporal and 12.7 cm spatial resolution along the attached cable. Note that in double-ended configuration both ends are measured sequentially and hence after processing the temperature observations were available with 2.3 s resolution.
Two calibration locations were created next to the DTS instrument and both ends of the cable were guided through them providing altogether four calibration sections in the data. The calibration sections were at minimum 10 meters long. Cold calibration point was created by digging the cable to 0.7 m depth in the peat soil. This depth was below the soil water level and hence the temperatures at that depth did not vary rapidly. Typically, the temperatures were below the air temperature and varied between 2.9 $^{\circ}C$ and 10.7 $^{\circ}C$ during the campaign. Warm calibration point was created with 50-l calibration bath filled with water and it was located in the temperature controlled cabin. Water heater was installed on June 28 and it was set to keep the water at 30 C. Prior to that the water temperature followed indoor air temperature in the temperature controlled cabin. Temperatures at the calibration locations were monitored with reference thermometers (PT-100) supplied with the DTS instrument.
[…]
Due to heterogeneity in the surface conditions (see Sect. 22.1), the cable height above the soil changed from one location to another. To avoid any terrain-induced curvature effects on the cable orientation, we decided not to follow the local terrain but to fasten the cable to wooden support structures at a constant height above sea level. This height was identified at each location with a handheld GNSS device (Geo 7X, Trimble Inc., USA), utilized to position the support structures in the area. At the location of the eddy covariance (EC) station, the cable height was around 3.1 m above the soil (see Fig. 1b). However, the local DTS measurement height varied and was approximately between 2 to 4 m above the soil and between 0.5 to 2.5 m above the local roughness elements (e.g. vegetation). Moreover, occasionally the cable was also surrounded by the few isolated trees left at the clearcut. For practical purposes, we ignore these local height variations and consider the height of the DTS cable to be 3.1 m above the soil in order to be consistent with the EC measurements.
To quantify the turbulent fluxes over such a complex surface, the EC equipment consisted of a 3-D sonic anemometer (uSonic-3 Cage MP, METEK GmbH, Germany) and an Infrared gas analyzer (LI-7200RS, LI-COR Biosciences, NE, USA). These together measured the three wind velocity components, sonic temperature, and carbon dioxide and water vapor mixing ratios at the height of 3.1 m above the soil. The infrared gas analyzer sampled the flow next to the sonic anemometer and all the EC data were logged with 10-Hz frequency. See more details on the EC measurement setup and flux footprint calculations in tikkasalo2025eddy.
2.3 DTS data processing
DTS data were post-field calibrated with Python (3.11.9) scripts and the processing followed des_tombe_estimation_2020, van_de_giesen_double-ended_2012 and peltola_probing_2022 with slight differences. Temperature can be estimated from the observed Stokes ($P_{s}$) and anti-Stokes ($P_{as}$) signals with the following equation (van_de_giesen_double-ended_2012):
$T_{x}(s,t)=\frac{\gamma}{\ln{\frac{P_{s}(s,t)}{P_{as}(s,t)}}+C(t)+\int_{0}^{s}\alpha(s^{\prime})ds^{\prime}},$ (1)
where $T_{x}=T_{x}(s,t)$ is the calibrated temperature at location $s$ along the fiber-optic cable at time step $t$, $\gamma$ and $C(t)$ are calibration parameters, and $\int_{0}^{s}\alpha(s^{\prime})ds^{\prime}$ describes the differential attenuation of $P_{s}$ and $P_{as}$ along the fiber. This integral was estimated from the forward and reverse signals for each location $s$ separately, corresponding to individual 30-min periods (van_de_giesen_double-ended_2012). Thereafter, at each time step, the calibration parameters $\gamma$ and $C$ were estimated for the forward and reverse signals by fitting Eq. 1 to $P_{s}$ and $P_{as}$ observed with the DTS instrument and $T$ measured with the reference sensors. Although, in theory, $\gamma$ should be constant, here it was treated as a fitting parameter varying in time. These parameters were used to calculate separate estimates of $T_{x}(s,t)$ from the forward ($T_{f}$) and reverse ($T_{r}$) signals. Since $T_{f}$ and $T_{r}$ were measured sequentially, $T_{r}$ data were linearly interpolated to match $T_{f}$ time stamps.
Noise estimates ($\sigma_{T}$, 1-$\sigma$ variability related to random noise) of $T_{f}$ and $T_{r}$ observations were computed for each 30-min period as,
$\sigma_{T}^{2}(s)=\frac{\overline{\left(T_{x}(s,t)-T_{x}(s+\Delta s,t)\right)^{2}}}{2},$ (2)
where $T_{x}$ represents either $T_{f}$ or $T_{r}$, $\overline{()}$ denotes temporal averaging and $\Delta s$ is the spatial step along the cable (0.127 cm). In this equation, we assume that the variability between consecutive data points in space is dominated by noise. The calibrated temperature $T$ at each $t$ and $s$ was calculated as a weighted average of $T_{f}$ and $T_{r}$ where inverse of noise estimates were used as weights (des_tombe_estimation_2020). Then the noise estimates for $T$ were calculated following noise propagation of weighted averaging. The spatial resolution of the data were reduced to 30 cm with spatial aggregation using weighted averaging as above and noise estimates for these spatial aggregates were calculated using noise propagation as above.
After this calibration procedure, we obtained calibrated $T$ at each time step $t$ and location $s$ along the cable. The $T$ data measured along the cable were converted to physical coordinates (ETRS89/TM35FIN coordinate system) using the measured cable locations (Sect. 2.2). Since there were double cabling at each physical location, the double measurements made at each location were further averaged using weighted averaging as above with the noise estimates obtained from the procedure described above. By utilizing this procedure, we obtained calibrated $T$ data with 2.3 s temporal and 30 cm spatial resolution throughout the measurement campaign at well-defined physical ($x$,$y$,$z$) locations (red lines in Fig. 1) over the study area. These calibrated temperature values were compared against the sonic anemometer measurements at the clearcut showing satisfactory agreement between the two (not shown).
2.4 Convection speed estimation
Figure 2: (a) The space-time contour plot of temperature fluctuations ($T^{\prime}$) is shown from the DTS measurements of a specific section (Section-1) for a particular 30-min period between 10:30-11:$0$ UTC on 18 May 2024. The spatial locations along the cable section are denoted by $X$ and the temperature fluctuations are obtained by linearly detrending the temperature signals measured along time. The white horizontal strips indicate the regions where the DTS cable was fastened to masts and the temperature values were omitted from those locations. (b) The space-time correlation curves of temperature fluctuations ($R_{TT}(\tau,r)$) are shown for the same 30-min period obtained from Section-1. The temporal lags or leads are denoted as $\tau$ while the colors indicate different values of spatial separations along $X$, $r$. The blue squares denote the peaks of the correlation curves ($\tau_{p}$) for each value of $r$. (c) A scatter plot between $r$ and $\tau_{p}$ is shown. The thick black color denotes the best fit line to compute the convective speed of the temperature structures, denoted as $U_{C}$.
To use the DTS and EC datasets, we focus on buoyancy-dominated conditions when the temperature signals were sufficiently strong (see Fig. 10 of Appendix A for further discussion). In particular, we selected those 30-min periods when there was no rain, incoming shortwave radiation was larger than 300 $\rm W\ m^{-2}$ and it was not strongly varying within the 30-min periods, the sensible heat flux ($50\rm\ Wm^{-2}$) and variance of temperature ($0.1\rm\ K^{2}$) were sufficiently large, air temperature did not exhibit strong trends, and the direction of the mean wind was within $\pm 20^{\circ}$ of the DTS cable positions (see Fig. 1b). This wind direction criterion resulted in two different scenarios. For Sections 1 to 5, the mean wind was parallel to the forest edge while for Section 6 it was perpendicular to the edge (see Fig. 1b). After applying all these conditions, we ended up with a total of 35 hours of DTS and EC data, or in other words, 70 half-hour periods. The DTS dataset was used to estimate the convective speeds of the turbulent structures by using the temperature signal as a tracer. The EC dataset served as a testbed to evaluate the impact of these convective speeds on the routine flux measurements.
Out of all our selected runs, for a specific representative 30-min period (10:30-11:00 UTC, 18 May 2024), we show the space-time contour plot of temperature fluctuations from Section 1 of the DTS cable (Fig. 2a). The temperature values were omitted from those locations where the DTS cable was fastened to wooden structures. This decision was undertaken to reduce any positive bias in the temperature signal since these wooden structures were not entirely thermally inert. Corresponding to each cable section, these omitted values represented around 6% of the cable length. Nevertheless, to ensure continuity in the data, those missing values were replaced with linearly-interpolated ones. Thereafter, the temperature fluctuations ($T^{\prime}$) were computed by linearly detrending the data over time separately at each location along the cable sections. The distance along Section 1 is denoted as $X$. To compute the convective speed of the temperature structures, we construct the space-time correlation curves from the DTS dataset. Mathematically, the two-point correlation functions of temperature fluctuations can be defined as,
$R_{TT}(\tau,r)=\frac{\overline{T^{\prime}(X,t)T^{\prime}(X+r,t+\tau)}}{\sigma_{T(X)}\sigma_{T(X+r)}},$ (3)
where $\tau$ are the temporal lags or leads, $r$ denotes the spatial separations along $X$, and $\sigma_{T(X)}$, $\sigma_{T(X+r)}$ are the standard deviations of temporal temperature fluctuations at locations $X$ and $X+r$. For the same data as shown in Fig. 2a, the quantity $R_{TT}(\tau,r)$ are plotted in Fig. 2b. Here, different colors represent the increasing values of $r$ while the blue squares indicate the peaks ($\tau_{p}$) attained by $R_{TT}(\tau,r)$ for a specific $r$. By fitting a straight line to a scatter plot between $r$ and $\tau_{p}$, the bulk convective speed of the temperature structures can be estimated (see the solid black line in Fig. 2c). The goodness of the fit can be adjudged by its $R^{2}$ value, which was well-beyond 0.9 in this example case (Fig. 2c). This linear relationship between $\tau_{p}$ and $r$ can be described as, $r=U_{C}\tau_{p}$, where $U_{C}$ is the bulk convective speed ($\rm m\ s^{-1}$). We restrict the $r$ values within 21 m while making the scatter plot, since for $r>21\rm\ m$, the $R_{TT}(\tau,r)$ values are small and therefore the estimations of $\tau_{p}$ are not statistically robust. To test the robustness of this method for atmospheric surface layer flows, we repeated this analysis on the cheng2017failure dataset (see Fig. S1 of the Supplementary Material). Similar to Figs. 2b–c, in Fig. S1 (Supplementary Material), the peak positions of $R_{TT}(\tau,r)$ curves shift towards increasing $\tau_{p}$ values as $r$ increases, thereby yielding a linear $r$-$\tau_{p}$ relationship. To ensure whether the shapes of $R_{TT}(\tau,r)$ curves were affected by the linear-interpolation operation, we increased the missing values to as large as 20% of the cable length and replotted Fig. 2b for this artificial data. We show this in Fig. S2 of the Supplementary Material and one could see that the peaks of $R_{TT}(\tau,r)$ curves behave analogously as in the original data.
Figure 3: (a) Contrary to Fig. 2b, the space-time correlation curves of temperature fluctuations ($R_{TT}(r,\tau)$) are shown where the $x$-axis represents the lags or leads in the spatial domain ($r$) while the colors indicate different values of temporal separations, $\tau$. The blue squares denote the peaks of the correlation curves ($r_{p}$) for each value of $\tau$. (b) A scatter plot between $r_{p}$ and $\tau$ is shown. The thick black color denotes the best fit line to compute the convective speed of the temperature structures, denoted as $U_{e}$.
Although the $r$-$\tau_{p}$ approach is more popular to compute the convective speed, everard2021sweeping argued that this method can give rise to high values due to the presence of sweeping effects associated with large turbulent eddies. Specifically, when the turbulence kinetic energy (TKE) of the flow is quite high, the sweeping effects have a strong domination on the estimates of $U_{C}$. To overcome this issue, everard2021sweeping suggested to use the $r_{p}$-$\tau$ relationship, which one can obtain by reversing the perspective, i.e. plotting the space-time correlation curves with respect to spatial lags or leads ($r$) at different values of temporal separations ($\tau$). For the same 30-min period akin to Fig. 2b, we plot the $R_{TT}(r,\tau)$ curves in Fig. 3a where the $x$-axis denotes the $r$ values while the colors of the curves indicate the values of $\tau$. The spatial peaks of $R_{TT}(r,\tau)$ curves are denoted as $r_{p}$ (see the blue squares in Fig. 3a). Similar to Fig. 2c, one can do a scatter plot between $r_{p}$ and $\tau$ (see Fig. 3b) to estimate a convection speed ($r_{p}=U_{e}\tau$), uncontaminated by the sweeping effects, which we denote as $U_{e}$. Upon comparing Fig. 2c with Fig. 3b, a difference emerges. In Fig. 3b, the $r_{p}$ values increase in a nearly consistent manner with $\tau$. On the other hand, in Fig. 2c, the $\tau_{p}$ values do not increase consistently but remain constant for certain ranges in $r$. This behavior occurs because the DTS data have a finer resolution in the spatial domain (30 cm) as compared to the temporal one (2.3 s).
After computing $U_{e}$, everard2021sweeping applied the elliptic model of he2010small to obtain the sweeping speed, $V$. The detailed derivation of the elliptic model of space-time correlation curves is provided in the Supplementary Material of everard2021sweeping, which we do not repeat here. By utilizing the relationship, $r=U_{C}\tau_{p}$, everard2021sweeping showed that the $V$ can be computed as (see their Eq. (5)),
$V=U_{e}\sqrt{\frac{U_{C}}{U_{e}}-1},$ (4)
where the non-zero values of $V$ describe any discrepancy between $U_{e}$ and $U_{C}$.
2.5 Scale-dependent convection speed
To compute the scale-dependent convection speeds of temperature structures, we follow the same procedure as outlined in cheng2017failure. By borrowing the concept of phase-dependent convection speeds from del2009estimation, cheng2017failure defined the convection speed at a wavenumber $\kappa$ ($U_{C}(\kappa)$) as,
$U_{C}(\kappa)=\frac{\Delta\varphi}{\kappa\times\Delta t}\$(5)
Here, $\Delta\varphi$ indicates the Fourier phase differences between the spatial temperature signals at time instants $t$ and $t+\Delta t$, $\kappa$ is the spatial wavenumber directly evaluated from the DTS measurements ($\kappa=2\pi/\lambda$, where $\lambda$ is the wavelength), and $\Delta t$ is the temporal lag (2.3 s for our DTS data). Unlike cheng2017failure, we used a wavelet based approach (torrence1998practical) to compute the phase differences between the spatial temperature signals at times $t$ and $t+\Delta t$, i.e. $\Delta\varphi$ . For our purposes, Morlet was used as the mother wavelet and upon comparison with a Fourier based approach, both yielded similar results (not shown).
3 Results and discussion
We begin with demonstrating the relationship between the convective speed estimations from the DTS measurements and the mean wind speed from the EC system at the clearcut. Corresponding to the EC data, $u$, $v$, and $w$ indicate the streamwise, cross-stream, and vertical velocities, while $T,q,c$ indicate temperature, moisture, and $\rm CO_{2}$ concentrations. After employing a double-coordinate rotation, the turbulent fluctuations in these quantities are computed through linear detrending. By using both DTS and EC datasets, the impact of scale-dependent convection speeds on the spectral properties of turbulence are discussed. We point out some potential limitations of the DTS measurements and how our findings influence in general the flux measurements from the eddy covariance systems.
3.1 Convection speeds of temperature structures
3.1.1 Bulk convection speeds
Figure 4: (a) The scatter plot between the convection speed estimated from the DTS measurements ($U_{C}$) and the mean wind speed from EC system ($U$) at the clearcut. The dash-dotted black line show the 1:1 relationship while the blue dash-dotted line is the best fit line. (b)–(c) The ratios $U_{C}/U$ are plotted against the turbulence intensities $\sigma_{u}/U$ for the cases when the direction of the mean wind was either along or perpendicular to the forest edge. For the perpendicular case, the wind directions are further subdivided when the wind was approaching directly from the forest or from the clearcut. (d) The scatter plot between $U_{e}$ and $U$ is shown, where $U_{e}$ is obtained from the $r_{p}$-$\tau$ relationship. The black dash dotted line indicates the 1:1 relationship. (e) The relationship between the square of the sweeping speed ($V^{2}$) and turbulence kinetic energy (TKE) is shown for the DTS cable sections 1–5 (along-wind conditions). The TKE is obtained from the EC system in the middle of the clearcut. The black, magenta, and red dash-dotted lines indicate the relationships $V^{2}\approx\rm TKE$, $V^{2}\approx\rm 2.16\ TKE$, and $V^{2}\approx\rm 2.25\ TKE$, respectively. The latter two are obtained from everard2021sweeping for flows over a vineyard canopy.
Figure 4a shows the scatter plot between the convective speed of temperature structures ($U_{C}$) estimated from the space-time correlation curves (Fig. 2b–c) and the mean wind speed from the EC system ($U$) at the clear cut. The gray-colored squares denote the Sections 1 to 5, and these correspond to the scenario when the wind was blowing parallel to the forest edge (along-wind). The red-colored squares denote Section 6 and correspond to the scenario when the direction of the mean wind was perpendicular to the forest edge (perpendicular-wind). The black dash-dotted line indicates 1:1 agreement between $U_{C}$ and $U$. For the along-wind cases, it is clear that the $U_{C}$ values are consistently larger than $U$, albeit with some scatter (gray-colored squares in Fig. 4a). Therefore, the temperature structures are advected faster than the local mean wind speed and the relationship between $U_{C}$ and $U$ can be expressed as, $U_{C}=0.77U+1.04$ (see the blue dash-dotted line in Fig. 4a). The same relationship describes the perpendicular-wind cases as well (red-colored squares), but the scatter is significantly increased.
According to the study by willis1976use, the Taylor’s hypothesis is rendered invalid when the turbulent intensities of the streamwise velocities ($\sigma_{u}/U$, where $\sigma_{u}$ is the standard deviation) exceed 0.5. In Figs. 4b–c, we show the scatter plots between $U_{C}/U$ and $\sigma_{u}/U$. Due to the scarcity in the number of samples, we cannot verify whether $U_{C}/U$ remains close to 1 when $\sigma_{u}/U\leq 0.5$. However, for values other than 0.5, the relationship between $U_{C}/U$ and $\sigma_{u}/U$ is linear with an asymptote of 1.9 as $\sigma_{u}/U>0.8$. All these relationships can be condensed to a singular form,
$U_{C}/U=(3\sigma_{u}/U-0.5)[\mathcal{H}(\sigma_{u}/U-0.5)-\mathcal{H}(\sigma_{u}/U-0.8)]+1.9\mathcal{H}(\sigma_{u}/U-0.8),$ (6)
where $\mathcal{H}(x)$ is a Heaviside function with $\mathcal{H}(x)=0$ when $x\leq 0$ and 1 otherwise.
For the perpendicular-wind cases, the mean wind directions can be further separated into two categories; when the wind was blowing directly from the forest and when it approached the forest edge from the clearcut. For the first category (forest), it appears that the ratios $U_{C}/U$ were closer to unity despite the changes in $\sigma_{u}/U$ (Fig. 4c). However, for the latter category (clearcut), the pink squares in Fig. 4c followed Eq. 6 more closely. This difference is not trivial by considering the fact that when the wind blows from the forest, the turbulence at the clearcut location has not fully equilibrated to the local surface conditions. On the other hand, when the wind approaches the forest edge from the clearcut, the situation reverses.
To verify this argument, in Fig. S3 of the Supplementary Material, we show the premultiplied vertical velocity spectra ($S_{ww}(f), f=f$= frequency) from the sonic anemometer observations at locations just above the forest, at horizontal distances of one and two canopy heights ($1h$ and $2h$, where $h=$ canopy height) from the forest edge, and at the clearcut. Except the above forest one, the spectra from the rest of these locations behave very similarly for the along-wind cases (Fig. S3c) and for the cases when the wind approached the forest edge from the clearcut (Fig. S3b). In particular, under the surface layer scaling, the spectra at the clearcut and at distances one and two canopy heights collapsed reasonably well. This is not true for the wind direction from the forest edge, where the spectra at all these locations showed a poor collapse under the same scaling and remained quite different from one another (Fig. S3a). Upon a closer look at Fig. S3a, corresponding to lower frequencies, $fS_{ww}(f)$ values at a distance of $1h$ were comparable to the above forest ones, and gradually decreased with increasing distances from the forest edge.
3.1.2 Influence of sweeping effects
From Figs. 4b–c, one important aspect to consider is that both the along- and perpendicular-wind scenarios are highly turbulent, with $\sigma_{u}/U$ values exceeding 0.5. flesch_wind_1999 observed highly turbulent wind conditions in a forest clearcut approximately at a distance of three to five times the canopy height. This distance matches with the location of our EC set up, situated approximately five times the canopy height from the forest edge. In such highly turbulent conditions, for along-wind cases and when the direction of the mean wind is from the clearcut location, $U_{C}$ values, corresponding to the range $0.5<\sigma_{u}/U<0.8$, increase with $\sigma_{u}/U$. everard2021sweeping pointed out that in highly turbulent conditions, the estimations of $U_{C}$ are influenced by sweeping effects associated with large eddies of the flow. They postulated that the occurrences of the sweeping events increase considerably when the TKE of the flow is quite high. To obtain an estimate of the convective speeds, uncontaminated by the sweeping effects, they suggested to use $U_{e}$ instead (see Fig. 3).
For our clearcut conditions, in Fig. 4d, we show a scatter plot between $U_{e}$ and $U$. One immediately recognizes two distinct features. First, contrary to Fig. 4a, in Fig. 4d, a majority of data points lie below the 1:1 line, thereby suggesting that the $U_{e}$ values remain mostly smaller than both $U$ and $U_{C}$. This is exactly the same conclusion reached by everard2021sweeping for their flow configuration over a vineyard canopy. Second, the scatter among the points in Fig. 4d remains considerably larger than in Fig. 4a. This is a similar feature observed by everard2021sweeping for their vineyard canopy dataset, where they found the relationship between $U_{e}$ and $U$ did not exactly follow the 1:1 line (see their Fig. 3a). However, within the limit of experimental accuracy, they ignored this scatter and considered $U_{e}\approx U$. Specific to our set up, for the along-wind conditions (Sections 1 to 5), after ensemble averaging the results in Fig. 4d, one could find that $U_{e}$ remained approximately equal to $U$. On the other hand, for the perpendicular-wind cases (shown by red squares), the number of data points in Fig. 4d is only four. This is because, nearly for all the perpendicular-wind periods, the $r_{p}$-$\tau$ relationships could not described as a best-fit straight line with enough statistical accuracy. The statistical accuracy is quantified in terms of whether the $R^{2}$ values of the linear fits were greater than 0.9. Since the computation of $U_{e}$ is tied to the elliptic approximation model of he2010small, we hypothesize that the accuracy of this model to describe the space-time correlation curves is impacted in situations when the presence of an upstream forest canopy disturbs the flow over a clearcut location. This limitation forced us to compute the sweeping speeds ($V$) for the along-wind conditions only.
everard2021sweeping showed that the square of the sweeping speeds ($V^{2}$) was intimately connected to the TKE of the flow. They found this relationship to be $V^{2}\approx\rm 2.16\ TKE$ for heights above a vineyard canopy and $V^{2}\approx\rm 2.25\ TKE$ for heights within the canopy (see the red and magenta dash-dotted lines in Fig. 4e). To test their prediction, in Fig. 4e, we show the scatter plot between $V^{2}$ and TKE for along-wind cases. Note that the TKE is independently estimated from the EC system at the middle of a clearcut, defined as $(\sigma_{u}^{2}+\sigma_{v}^{2}+\sigma_{w}^{2})/2$. Similar to everard2021sweeping, we do see a strong relationship between the two. However, as opposed to everard2021sweeping, the relationship in Fig. 4e emerges to be $V^{2}\approx\rm TKE$ (black dash-dotted line in ig. 4e). This discrepancy is consistent with the expectation that the boundary conditions on the flow could potentially alter the relationship between $V^{2}$ and TKE (everard2021sweeping). Notwithstanding this difference, in agreement with everard2021sweeping, one can conclude that in highly turbulent conditions, the sweeping effects indeed become quite important and cause the $U_{C}$ values to appear significantly larger than the mean wind speed. The sweeping effects are also associated with modifying the turbulence frequency spectrum (everard2021sweeping), which will become apparent when we discuss the temporal to spatial conversion of turbulence spectra using scale-dependent convection speeds (see Sect. 3.2).
3.1.3 Scale-dependent convection speeds
Figure 5: (a) The scale-dependent convection speeds ($U_{C}(\kappa z)$) are plotted against the streamwise wavenumbers ($\kappa$) computed directly from the DTS spatial information. The wavenumbers are normalized by the measurement height ($\kappa z$) while $U_{C}(\kappa z)$ values are scaled by the convection speed $U_{e}$. The gray- and red-colored lines denote the along- and perpendicular-wind cases, respectively. The error bars indicate the spread, computed as one standard deviation from the ensemble mean. (b) Same as in (a), but the $U_{C}(\kappa z)$ values are scaled by $U_{C}$. The various fits to the scale-dependent convection speeds are shown by the dash-dotted lines (see the legend).
Hitherto, we have discussed the bulk convective speeds but now we turn our attention towards the scale-dependent convection speeds. In Fig. 5a, we show the ensemble-averaged scale-dependent convection speed estimations from the spatial DTS data of temperature at all the six sections. As before, the Sections 1–5 correspond to along-wind cases while Section 6 represents the perpendicular ones. The convection speeds in Fig. 5a are normalized by the uncontaminated form of the bulk convective speed, $U_{e}$, and the wavenumbers ($\kappa$) computed from the spatial DTS separations are scaled with the measurement height ($z$). As $U_{e}$ values have significant run-to-run variations (Fig. 4d), we used an ensemble-averaged value of $U_{e}$ in Fig. 5a, separately for the six different sections. Specific to the perpendicular-wind section, this average was carried out only over four 30-minute periods (see the red squares in Fig. 4d). The choice of $z$ to scale $\kappa$ is in line with hilland2024systematic, who claimed that in the context of Taylor’s hypothesis the measurement height is indeed an important parameter. Ideally, one should use zdz-d, where $d$ is the displacement height, but here we used only $z$ since to estimate $d$ a detailed vertical profile of wind measurements is required and that was not available at our measurement site. This would not have any impact on the results as $d$ only acts to shift the $x$ coordinates horizontally.
Figure 5a depicts that at lower wavenumbers, $\kappa z<1$, $U_{C}(\kappa z)$ values tend towards nearly two to three times the values of $U_{e}$, irrespective of the wind directions. On the other hand, if one uses $U_{C}$ to scale the $U_{C}(\kappa z)$ values, at lower wavenumbers, $U_{C}(\kappa z)\to U_{C}$ (see Fig. 5b). At present we do not understand this difference and hence continue with normalizing the $U_{C}(\kappa z)$ values with $U_{C}$, despite acknowledging its limitations. At larger wavenumbers, $\kappa z>1$, $U_{C}(\kappa z)/U_{C}$ values decrease in a power-law fashion with the exponent being equal to ${(\kappa z)}^{-0.8}$ (Fig. 5b). This is surprisingly different from the result of cheng2017failure who reported this exponent to be ${(\kappa z)}^{-1/3}$. We replotted the results of cheng2017failure and indeed for different atmospheric stability conditions (near-neutral to highly-convective) their data agree more closely with ${(\kappa z)}^{-1/3}$ instead of ${(\kappa z)}^{-0.8}$ (see Fig. S4 in the Supplementary material). However, compared to ours, the data of cheng2017failure exhibited a significant amount of scatter. Before we discuss potential reasons behind the disagreement with cheng2017failure, it is important to recognize that, like any other observations, the DTS measurements are not perfect and are to a certain degree contaminated by noise and possibly also other instrument artifacts (peltola2021suitability). In Appendix B, we consider the impact of such noises on the 0.8-0.8 scaling exponent.
The apparent disagreement with the results of cheng2017failure could be attributed to two factors. First, the presence of anisotropy at smaller scales of the flow. cheng2017failure theoretically derived their 1/3-1/3 exponent for atmospheric surface layer flows by using the advection-diffusion equation and a power-law dependence in the turbulent diffusion coefficient at the inertial subrange scales. In fact, they considered the turbulent diffusion coefficient of the inertial subrange eddies to vary as $r^{4/3}$, where $r$ represent their sizes. This $4/3$ power-law is a direct consequence of the turbulent particle pair diffusion theory introduced by Richardson (richardson1926atmospheric). richardson1926atmospheric assumed that the turbulence is locally isotropic at the inertial subrange scales, and the energy spectrum follows a 5/3-5/3 slope. In recent times, malik2018turbulent has shown that in the limit of asymptotically infinite inertial subrange (symptomatic of atmospheric flows with very-high Reynolds number), the exponent of the power-law ($\gamma$) associated with turbulent diffusion can vary between $(1+p)/2<\gamma<2$, where $p$ is the slope of the energy spectrum in the inertial subrange. If $p$ is $5/3$, then this range turns out to be between $4/3<\gamma<2$.
While deriving this result, malik2018turbulent questions the assumption of locality by introducing a hypothesis that the non-local scales of motion do contribute to the particle pair diffusion process at inertial subrange scales. In the definition of malik2018turbulent, the sizes of these non-local eddies are larger than the inertial subrange ones but smaller than the largest scales, typically the integral scales of streamwise velocity fluctuations. Now, in our case, we obtain a power-law of 0.8-0.8. If one follows the same line of derivation as outlined in cheng2017failure, a 0.8-0.8 power-law can be obtained if the turbulent diffusion coefficient of the inertial subrange eddies could be expressed as, $r^{1.8}$. Although this range of $\gamma$ can justify a relationship of $r^{1.8}$ (since $4/3<1.8<2$), from point-based temporal measurements in atmospheric flows, it is nearly impossible to provide direct evidence of this non-locality. However, this can be done indirectly. The presence of non-locality at the inertial subrange scales impart anisotropy at those scales (katul1997energy). The results of chowdhuri2024quantifying have demonstrated that, for a wide range of both atmospheric and laboratory datasets, significant anisotropy can persist even at smaller scales of the flow. This is true for our EC dataset, as the flow remains anisotropic at the inertial subrange scales when computed through the intermittency-anisotropy framework of chowdhuri2024quantifying (not shown).
The second reason is related to turbulence intensities. For our heterogeneous flow set up (a forest clearcut), the conditions are highly turbulent with $\sigma_{u}/U>0.5$. cheng2017failure did not report the values of their turbulence intensities. Later on, han2022applicability reevaluated the dataset of cheng2017failure. They reported that the $\sigma_{u}/U$ values of the eight half-hour runs of cheng2017failure (spanning different stability conditions) were less than 0.5 with only one exception (see Table 1 of han2022applicability). Thus, cheng2017failure mostly encountered low turbulent wind conditions ($\sigma_{u}/U<0.5$). The study by everard2021sweeping has shown that in highly turbulent conditions, significant sweeping effects associated with large eddies can persist in the flow and can induce incorrect spectral scalings when converting the spectra from frequency to wavenumber domain by using only the mean wind speed. This indeed can be accounted for by using scale-dependent convection speeds. everard2021sweeping established this by considering a limiting case with zero-mean wind speed but with non-zero sweeping effects (see their Eqs. 6 and 9). Therefore, apart from anisotropic effects at smaller scales, the bulk turbulence intensities could also have its impact on the exponents of scale-dependent convection speeds. Next, we provide a recipe to convert the spectra from temporal to spatial domain by using scale-dependent convection speeds.
3.2 Conversion of turbulence spectra
3.2.1 Frequency to wavenumber conversion
Under the assumption that the turbulent structures of different scales advect past the measurement location at similar speeds, one could essentially transform the frequencies ($f$) to wavenumbers ($\kappa$) using $\kappa=2\pi f/U$, where $U$ is the mean wind speed estimated from the sonic anemometer. Irrespective of whether one uses $U$, $U_{C}$, or $U_{e}$ to convert the frequencies to wavenumbers, it is a linear transformation. However, if the convective speeds of turbulent structures are scale-dependent, it introduces additional complexities to the transformation. From our DTS observations, it is evident that $U_{C}(\kappa z)$ is equal to $U_{C}$ at scales $\kappa z<1$, while it varies as ${(\kappa z)}^{-0.8}$ at scales $\kappa z>1$ (see Fig. 5b). In general, one could express this dependence through the following functional form,
$\frac{U_{c}(\kappa z)}{U_{C}}=\frac{1}{{(1+\kappa z)}^{\alpha}},$ (7)
with $0\leq\alpha<1$. Here, $\alpha=1$ corresponds to a limit for the randomly-shuffled samples (see Fig. 11a) while $\alpha=0$ indicates no scale-dependency. If $\alpha=0.8$, the functional form of Eq. 7 fits the data exceptionally well (see the yellow dash-dotted line in Fig. 5b). Now, if this relationship is used to convert the frequencies to wavenumbers, one could write,
$\kappa z=\frac{2\pi fz}{U_{C}(\kappa z)}.$ (8)
By replacing Eq. 7 in Eq. 8, one could rewrite Eq. 8 as,
$\kappa z=\frac{2\pi fz}{U_{C}}{(1+\kappa z)}^{\alpha}.$ (9)
Depending on $\alpha$, Eq. 9 is a $\rm n$-th order polynomial of $\kappa z$. For instance, when $\alpha=0.8=4/5$, Eq. 9 becomes,
${(\kappa z)}^{5}=a{(1+\kappa z)}^{4},$ (10)
where $a=(2\pi fz/U_{C})^{5}$. By solving the polynomial in Eq. 9 and restricting its roots to positive values, one could convert the frequencies to wavenumbers by using a scale-dependent convection speed.
It is important to notice that since Eq. 9 is a non-linear function, the wavenumbers generated from Eq. 9 will not be spaced equally in a linear space. Accordingly, the wavenumber increments ($d\kappa$) themselves would depend on the values of $\kappa$. In fact, by differentiating Eq. 9, one can show,
$d(\kappa z)=\frac{{(1+\kappa z)}^{\alpha+1}}{1+(1-\alpha)\kappa z}d(\frac{2\pi fz}{U_{C}}).$ (11)
We next discuss the implications of Eqs. 9 and 11 on spectral conversion.
3.2.2 Spectral conversion
The spectral energy at a specific wavenumber $\kappa z$ ($E_{xx}(\kappa z)$, where $x$ could be either velocity or scalar) is evaluated by integrating the spectral energy density between the limits $\kappa z$ and $\kappa z+d(\kappa z)$, and expressed as (within a specific constant),
$E_{xx}(\kappa z)\sim S_{xx}(\kappa z)d(\kappa z),$ (12)
where $S_{xx}(\kappa z)$ is the spectral energy density. In the absence of spatial information, one could apply Taylor’s hypothesis to convert the frequencies to wavenumbers. In that case, one could write Eq. 12 as,
$E_{xx}(\kappa z)\sim S_{xx}(2\pi fz/U_{c})d(2\pi fz/U_{c}).$ (13)
Let us consider that in the inertial subrange scales ($\kappa z\gg 1$) the spectral energy density in the temporal domain vary as $f^{-\beta}$ while at larger scales ($\kappa z\ll 1$) it varies as $f^{-\gamma}$. Therefore, under such conditions, one could write Eq. 13 as,
$E_{xx}(\kappa z)\sim\begin{cases}{(2\pi fz/U_{c})}^{-\gamma}d(2\pi fz/U_{c})&{\kappa z\ll 1}\\ {(2\pi fz/U_{c})}^{-\beta}d(2\pi fz/U_{c})&{\kappa z\gg 1}\end{cases}$ (14)
The Taylor’s hypothesis is valid when $\alpha=0$. However, when $\alpha$ is non-zero, one needs to replace $2\pi fz/U_{c}$ and $d(2\pi fz/U_{c})$ with Eqs. 9 and 11, respectively. By doing such replacements, one ends up with,
$\begin{split}E_{xx}(\kappa z)&\sim{\left[\frac{\kappa z}{{(1+\kappa z)}^{\alpha}}\right]}^{-\gamma}\left[\frac{1+(1-\alpha)\kappa z}{{(1+\kappa z)}^{\alpha+1}}\right]{d(\kappa z)}\\ E_{xx}(\kappa z)&\sim{\left[\frac{\kappa z}{(\kappa z)^{\alpha}(1+\frac{1}{\kappa z})^{\alpha}}\right]}^{-\beta}\left[\frac{1+(1-\alpha)\kappa z}{{(1+\kappa z)}^{\alpha+1}}\right]{d(\kappa z)}\end{split}$ (15)
When $\kappa z\ll 1$,
$\begin{split}{(1+\kappa z)}^{\alpha}&\to 1\\ \frac{1+(1-\alpha)\kappa z}{{(1+\kappa z)}^{\alpha+1}}&\to 1.\end{split}$ (16)
However, at inertial subrange scales, $\kappa z\gg 1$, and hence one could simplify the terms in Eq. 15 as,
$\begin{split}\left(1+\frac{1}{\kappa z}\right)^{\alpha}&\to 1\\ {(1+\kappa z)}^{\alpha+1}&={(\kappa z)}^{\alpha+1}\left(1+\frac{1}{\kappa z}\right)^{\alpha+1}\to{(\kappa z)}^{\alpha+1}\\ [1+(1-\alpha)\kappa z]&\to(1-\alpha)\kappa z\end{split}$ (17)
By inserting these approximations of Eqs. 16 and 17 in Eq. 15, one obtains,
$E_{xx}(\kappa z)\sim\begin{cases}{(\kappa z)}^{-\gamma}d(\kappa z)&{\kappa z\ll 1}\\ (1-\alpha){(\kappa z)}^{-\beta(1-\alpha)}(\kappa z)^{-\alpha}d(\kappa z)&{\kappa z\gg 1}\end{cases}$ (18)
For $\kappa z\gg 1$, Eq. 18 could be further rearranged as,
$E_{xx}(\kappa z)\sim(1-\alpha){(\kappa z)}^{-\beta+(\beta-1)\alpha}d(\kappa z).$ (19)
This is the desired result and one could see that the scale-dependent conversion does not change the spectral slope at smaller wavenumbers ($\kappa z\ll 1$) but yields a spectral slope of $\beta+(\beta-1)\alpha$ in the inertial subrange ($\kappa z\gg 1$) when the temporal frequencies are converted to wavenumbers using Eq. 7. The spectral behaviors described by Eq. 18 are valid for low- and high-wavenumber ranges. For the mid-wavenumber range, additional terms will appear after carrying out the binomial expansions of ${(1+\kappa z)}^{\alpha}$ and ${(1+\kappa z)}^{\alpha+1}$. Specific to the inertial subrange ($\kappa z\gg 1$), if $\beta>1$, the extra factor of $(\beta-1)\alpha$ in Eq. 19 implies that applying Taylor’s hypothesis underestimates the turbulent spectrum in the inertial subrange when $\alpha$ is non-zero. By comparing Eq. 19 with the findings of cheng2017failure, they obtained $\beta=2$ and $\alpha=1/3$ in the inertial subrange scales (see their Figs. 3 and 5). After replacing these values in Eq. 19, one obtains $E_{xx}(\kappa z)\sim{(\kappa z)}^{-5/3}d(\kappa z)$, which is precisely the result of cheng2017failure (see Section 2 in the Supplementary Material where we establish this equivalence). For our flow set up, $\beta$ appears to be $5/3$ (see Fig. 12a), and therefore the conversion yields, $5/3+2\alpha/3$. Note that a similar conversion can be carried out for the flux cospectra, with only difference being the slope $\beta$ will have a different value than the spectra (wyngaard1972cospectral). An important aspect to note is for the values of $0\leq\alpha<1$ and $\beta\geq 1$, the upper bound of the spectral slope in Eq. 19 is always 1-1. In essence, by applying Eqs. 9 and 11 on any temporal turbulence dataset, one could convert the spectra from temporal to spatial domain at any specific value of $\alpha$.
3.2.3 DTS temperature spectra
Figure 6: Corresponding to section-1 of the DTS cable, (a) the original and converted DTS temporal spectra are displayed. The original DTS temporal spectrum is shown as blue lines with square markers, while the converted one using scale-dependent convection speeds is shown as a black line with square markers. The conversion yields a spectral slope of $5/3+0.8\times 2/3\approx-1.13$ in the temporal domain (see Eq. 19), illustrated as a magenta line. (b) The spatial DTS spectrum is shown, and the magenta line there indicates the same slope as obtained in the temporal domain after conversion.
To test whether this procedure generates meaningful results, we used the DTS dataset of temperature, since that encompasses both temporal and spatial information. For illustration purposes, corresponding to Section-1 of the DTS cable, in Fig. 6, we show the temperature spectra for both the temporal (Fig. 6a) and spatial (Fig. 6b) domains. For the spatial spectra in Fig. 6b, the wavenumbers were computed directly from the DTS spatial separations by using $\kappa=2\pi/\lambda$, where $\lambda$ is the wavelength. In Fig. 6a, the blue lines with square markers indicate the spectra of temperature fluctuations in a temporal domain, where the frequencies were converted to wavenumbers ($\kappa=2\pi f/U_{C}$) using the bulk convective speed $U_{C}$. The temporal spectra in Fig. 6a were replotted with scale-dependent convective speeds, where Eq. 10 was used to convert the frequencies to wavenumbers while the spectral densities were calculated using Eq. 11 (black lines with square markers). Mathematically, this conversion can be expressed as,
$\frac{E_{TT}(2\pi fz/U_{c})}{d(2\pi fz/U_{c})}\equiv\frac{{E_{TT}\left(\frac{\kappa z}{{(1+\kappa z)}^{\alpha}}\right)}\left(1+(1-\alpha)\kappa z\right)}{{{(1+\kappa z)}^{\alpha+1}}d(\kappa z)},$ (20)
where $E_{TT}$ is the spectral energy, computed as the squared magnitudes of the complex Fourier coefficients of temperature. By using $\alpha=0.8$ in Eq. 20 and applying Eq. 18, one recovers a converted temporal spectra with a slope of $5/3+(5/3-1)\times 0.8\approx-1.13$ in the inertial subrange scales (see the magenta line in Fig. 6a). Upon comparison with Fig. 6b, it is clear that the spectra in the spatial domain too show this enhanced slope at larger wavenumbers (see the magenta line in Fig. 6b).
Figure 7: (a)–(f) The converted DTS temporal spectra and spatial spectra are compared for the DTS cable sections, spanning from 1 to 6. Sections 1-5 describe the along-wind periods (a–e) while Section 6 (f) describes the perpendicular-wind periods. The black dash dotted lines in (a)–(f) indicate the 5/3-5/3 spectral slope.
In Figs. 7a–f, we show the comparisons between the converted temporal spectra (gray squares) and spatial spectra (red squares) for all the six sections of the DTS cable. As noted before, Sections 1-5 describe the along-wind periods (Figs. 7a–e) while Section 6 describes the perpendicular-wind periods (Fig. 7f). The error bars denote one standard deviation from the ensemble mean. Notwithstanding the sections, the spectra in temporal space agree well with the spatial spectra, when the frequencies are converted to wavenumbers using Eqs. 10 and 11. While carrying out such conversion, we use $\alpha=0.8$ in Eq. 11. After establishing this, we do the same exercise with the EC measurements. However, instead of restricting ourselves to a specific value of $\alpha$, we carry out the spectral conversion on EC dataset for a range of $\alpha$ values. This is done to systematically investigate the impact of scale-dependent conversion on the EC measurements of turbulence statistics.
3.3 Impact on EC measurements
3.3.1 EC spectra
Figure 8: For the along-wind periods, the ensemble averaged EC spectra of (a) $u$, (b) $w$, (c) $T$, and (d) heat cospectra (wTwT) are shown. The different colored lines indicate the temporal spectra being converted according to a specific exponent of scale-dependent convection speeds ($\alpha$), which vary from 0 to $0.8$ at increments of $0.1$. Here $\alpha=0$ indicates that the conversion has been achieved through the Taylor’s hypothesis. The black lines in (a)–(c) indicate the spectral slope of 5/3-5/3 while the red line indicates a slope of $5/3+0.8\times 2/3\approx-1.13$ (see Eq. 19). For (d), the black line indicates a 7/3-7/3 cospectral slope while the red line indicates a slope of $7/3+0.8\times 4/3\approx-1.3$ (see Eq. 19). For the same along-wind component, in (e)–(h), the flux cospectra of momentum ($uw$), heat ($wT$), moisture ($wq$), and $\rm CO_{2}$ (wcwc) are shown in premultiplied form. Similar to (a)–(d), the different colored lines indicate the range of $\alpha$ values. For all the panels, the spectral and cospectral amplitudes are normalized by respective variances and the product of the standard deviations.
In Figs. 8a–c, we show the ensemble averaged $u$, $w$, and $T$ spectra from the EC dataset, corresponding to the periods when the wind was blowing parallel to the forest edge (along-wind scenario). The conclusions remain the same for the perpendicular-wind periods too, and therefore not being shown. The different colored lines in Figs. 8a–c represent the spectra being converted from temporal to spatial domain using scale-dependent convection speeds (see the legend in Fig. 8a). In particular, we use the following relationship to describe the scale-dependent convection speeds, $U_{C}(\kappa z)/U_{C}=1/(1+\kappa z)^{\alpha}$, with $0\leq\alpha<1$. We choose a range of $\alpha$ values varying between 0 to 0.8 at an increment of 0.1. These conversions are carried out following the Eqs. 9 and 11. Note that $\alpha=0$ indicates the situation with no scale-dependency and hence equivalent to what one would obtain by applying Taylor’s hypothesis. While conducting this analysis, we do consider that the scale-dependent convection speeds have a power-law relationship at smaller scales of the flow. Accordingly, these spectral conversions are sensitive to $\alpha$ values rather than whether $U$, $U_{C}$, or $U_{e}$ were used to scale $U_{C}(\kappa z)$.
At a first glance, these spectral results in Figs. 8a–c appear counter-intuitive. This is because, the temporal spectra, without any application of scale-dependent conversion, follow a clear 5/3-5/3 slope (see the black dash-dotted lines in Figs. 8a–c). In accordance with Eq. 19, any non-zero $\alpha$ values would yield a spectral slope of $\beta+(\beta-1)\alpha$ in the inertial subrange scales, where $\beta$ (=5/3=5/3 for our case) is the slope in the frequency domain ($\alpha=0$). Since $5/3>1$, the term $(\beta-1)\alpha>0$, and hence the spectral slopes will exceed 5/3-5/3 when the scale-dependent conversion is carried out, irrespective of whether $\alpha$ is $1/3$ or $0.8$. This is precisely what we observe in Figs. 8a–c that the spectra become progressively flattened as $\alpha$ increases. In that respect, we agree with cheng2017failure that applying Taylor’s hypothesis systematically underestimates the turbulent spectrum in the inertial subrange. Specific to $\alpha=0.8$, which is the value we obtain for our dataset (see Fig. 5b), the converted spectral slopes are equal to $5/3+0.8\times 2/3\approx-1.13$ (see the red dash-dotted lines in Figs. 8a–c).
It is prudent to ask, what does it physically mean for the converted temporal spectra, at an $\alpha$ value of 0.8, to show a broad region of $\approx-1.13$ spectral slope with no 5/3-5/3 region?. As shown in Figs. 6b and 7, the spatial spectra of DTS temperature fluctuations show a similar enhanced slope ($\approx-1.13$) at the inertial subrange wavenumbers ($\kappa z\geq 1$). Since the DTS dataset is limited by temperature information only, we do not have equivalent spatial information for the velocity counterpart. In the present scenario, by assuming that the turbulence structures in the velocity fields follow the same scale-dependent convection speed relationship as of temperature, we consider that the same spectral enhancement at the inertial subrange scales occur for $u$ and $w$ components. This of course needs to be validated through other means than DTS measurements (perhaps large eddy simulations) but that is beyond the scope of this work. Keeping such limitation in mind, we explore two potential reasons for this spectral enhancement.
3.3.2 Dimensional analysis
The first reason is small-scale intermittency that could cause deviations from the expected 5/3-5/3 spectral slope (dubrulle1994intermittency). However, such corrections are small for the second-order turbulence statistics (i.e. spectra) and therefore cannot explain our observations. The second reason is turbulence anisotropy imparted by the large-scale motions to smaller scales of the flow (chowdhuri2024quantifying). To account for such large‐scale inertial scale interactions, katul2025scaling proposed a dimensional analysis framework by considering an additional influence of large-scale field on the inertial subrange turbulence statistics (see their Eq. 15). Although katul2025scaling developed this framework for pressure fluctuations in near-neutral roughness sublayer flows, we extend the same to our case. To start with, let us consider that the turbulence velocity spectra in the inertial subrange scales depend on the large-scale velocity field, which we denote as $X$. For the scalar field, such as temperature, given the strong relationship between the sweeping speed of temperature structures and TKE of the flow (see Fig. 4e), we extend the same dependence on $X$ to the temperature spectra. Therefore, the spectral amplitudes of velocity and temperature in the inertial subrange scales have a functional dependence of the form,
$\begin{split}S_{xx}(\kappa)=F(\epsilon_{\rm TKE},\kappa,X)\\ S_{TT}(\kappa)=F(\epsilon_{\theta},\epsilon_{\rm TKE},\kappa,X).\end{split}$ (21)
In Eq. 21, $x=u,w$, $\epsilon_{\theta}$ and $\epsilon_{\rm TKE}$ are the dissipation rates of temperature variance and turbulence kinetic energy and $\kappa$ is the wavenumber. We now conduct a dimensional analysis and for that purpose, we consider the three fundamental dimensions $[L,T,K]$, where $L$, $T$, and $K$ represent the dimensions of length, time, and temperature, respectively. Since the dimensions of $S_{xx}(\kappa)$ and $S_{TT}(\kappa)$ are $L^{3}T^{-2}$ and $K^{2}L$, respectively, one could write Eq. 21 in its dimensional form as,
$\begin{split}L^{3}T^{-2}\sim\epsilon^{a_{1}}_{\rm TKE}\kappa^{b_{1}}X^{c_{1}}\\ K^{2}L\sim\epsilon^{a_{2}}_{\theta}\epsilon^{b_{2}}_{\rm TKE}\kappa^{c_{2}}X^{d_{2}},\end{split}$ (22)
where the dimensions of $\epsilon_{\theta}$, $\epsilon_{\rm TKE}$, $\kappa$, and $X$ are, $K^{2}T^{-1}$, $L^{2}T^{-3}$, $L^{-1}$, and $LT^{-1}$, respectively. Henceforth, we remove the subscript TKE from $\epsilon_{\rm TKE}$. By substituting these dimensions in Eq. 22, we get,
$\begin{split}L^{3}T^{-2}\sim L^{2a_{1}-b_{1}+c_{1}}T^{-(3a_{1}+c_{1})}\\ K^{2}L\sim K^{2a_{2}}L^{2b_{2}-c_{2}+d_{2}}T^{-(a_{2}+3b_{2}+d_{2})},\end{split}$ (23)
By equating the similar dimensions at both sides, and after some algebraic manipulations, one ends up with the following spectral forms,
$\begin{split}S_{xx}(\kappa)\sim{\epsilon}^{-(1+c_{1})}\kappa^{c_{1}}X^{5+3c_{1}}\\ S_{TT}(\kappa)\sim\epsilon_{\theta}{\epsilon}^{-(2+c_{2})}\kappa^{c_{2}}X^{5+3c_{2}}.\end{split}$ (24)
Now, in the place of $X$, we put the velocity scale ${(\rm TKE)}^{1/2}$, which represents the energetic large-scale motions, and further rewrite Eq. 24 as,
$\begin{split}S_{xx}(\kappa)\sim{\epsilon}^{-(1+c_{1})}\kappa^{c_{1}}{\rm TKE}^{(5+3c_{1})/2}\\ S_{TT}(\kappa)\sim\epsilon_{\theta}{\epsilon}^{-(2+c_{2})}\kappa^{c_{2}}{\rm TKE}^{(5+3c_{2})/2}.\end{split}$ (25)
If $c_{1}=c_{2}=-5/3$ in Eq. 25, one recovers the standard Kolmogorov spectra of temperature and velocity,
$\begin{split}S_{xx}(\kappa)\sim{\epsilon}^{2/3}\kappa^{-5/3}\\ S_{TT}(\kappa)\sim\epsilon_{\theta}{\epsilon}^{-1/3}\kappa^{-5/3},\end{split}$ (26)
with no dependence whatsoever on the velocity scale ${\rm TKE}^{1/2}$. However, if the spectral slopes are of the form $c_{1}=-5/3+\delta_{1}$ and $c_{2}=-5/3+\delta_{2}$, then Eq. 25 dictates,
$\begin{split}S_{xx}(\kappa)\sim{\epsilon}^{2/3-\delta_{1}}\kappa^{-5/3+\delta_{1}}{(\rm TKE)}^{3\delta_{1}/2}\\ S_{TT}(\kappa)\sim\epsilon_{\theta}{\epsilon}^{-(\delta_{2}+1/3)}\kappa^{-5/3+\delta_{2}}{(\rm TKE)}^{3\delta_{2}/2}.\end{split}$ (27)
The non-vanishing terms ${(\rm TKE)}^{3\delta_{1}/2}$ and ${(\rm TKE)}^{3\delta_{2}/2}$ in Eq. 27 imply that the presence of a large-scale velocity field can indeed cause deviations from the standard Kolmogorov spectral form. However, unlike the Kolmogorov spectra, in this case, the exponents $c_{1}$ and $c_{2}$ are not unique (hence not universal) and $c_{2}\approx-1.13$ emerges to be one of the possibilities. Apart from that, in order to satisfy Eq. 25, the exponents $c_{1}$ and $c_{2}$ need not be equal for the velocity and temperature signals. Therefore, by introducing the effects of the large-scale velocity field on inertial subrange scale motions positions us to have spectral exponents larger than 5/3-5/3 with differences being allowed to exist between the velocity and temperature signals. We clarify that our intent behind conducting this dimensional analysis is to provide a plausible reason for enhanced spectral slopes at inertial subrange scales, rather than deriving a new spectral model of velocity and temperature from first principles. For instance, it is currently unclear up to what wavenumber limit such enhancements in spectral slopes continue. Moreover, are the values of $c_{1}$ and $c_{2}$, satisfying Eq. 25, equal for velocity and temperature signals? To answer these questions more spatially explicit field measurements and numerical simulations are required, which is a matter of future research.
3.3.3 EC cospectra
Next, we investigate the impact of $\alpha$ on flux cospectra. In Fig. 8d we show the heat flux ($wT$) cospectra, normalized with the product of the standard deviations of vertical velocity ($\sigma_{w}$) and temperature ($\sigma_{T}$). Analogous to Figs. 8a–c, the heat flux cospectra have been transformed using Eqs. 9 and 11 with different $\alpha$ values. The $wT$ cospectral amplitudes, without any application of scale-dependent conversion, closely follow a 7/3-7/3 slope at larger wavenumbers (see the black dash-dotted line in Fig. 8d). As Eq. 19 predicts, this slope transforms to $7/3+(7/3-1)\times 0.8$ for $\alpha=0.8$. The exact cospectral slope has been debatable since their exponents have shown to be not unique and depend on how the dimensional analysis is performed (cheng2020power). Therefore, in that spirit, it is possible to incur a cospectral slope of $7/3+(7/3-1)\times 0.8$ in the spatial domain, which is nearly equal to 1.3-1.3 (see the red dash-dotted line in Fig. 8d). We extend this analysis to other fluxes, namely momentum, moisture and $\rm CO_{2}$.
It is natural to plot the cospectra in a premultiplied form since the area under this log-linear coordinate system is directly related to the flux contributions across different scales (kaimal1994atmospheric). In Figs. 8e–h, we show the premultiplied cospectra of momentum ($uw$), heat ($wT$), moisture ($wq$), and $\rm CO_{2}$ (wcwc) vertical fluxes. Similar to Figs. 8a–d, the different colored lines indicate different $\alpha$ values. As one would expect, the slopes of the premultiplied cospectra increase as $\alpha$ increases. Quite interestingly, the behavior of $uw$ (Fig. 8e) and $wc$ (Fig. 8h) cospectra remain different than heat (Fig. 8f) and moisture (Fig. 8g). It appears that the rate by which the cospectral slopes increase with $\alpha$ values remain smaller for momentum and $\rm CO_{2}$ fluxes. This occurs because, these results are ensemble averaged and the run-to-run variations in $uw$ and $wc$ cospectra are larger than what are present in the heat and moisture cospectra (not shown). Since we apply our spectral conversion technique on individual 30-min periods, the run-to-run variations could have impacted the results when they are ensemble averaged.
Notwithstanding such differences, it is clear that the scale-dependent conversion increases the slopes of cospectral curves. Particularly, the results from Figs. 8f and g demonstrate that non-zero $\alpha$ values can cause an appreciable increase in the heat and moisture fluxes at smaller scales of the flow. As conjectured by cheng2017failure, such enhancement could in fact increase the flux values when Eq. 19 is integrated across the entire inertial subrange scales and by doing so perhaps one could solve the enigma of surface energy imbalance (mauder2020surface). This remains a topic for our future research. However, as we show below, the frequency to wavenumber conversion through Eq. 9 yields a critical frequency limit beyond which the space-time conversion is uncertain.
3.3.4 A critical frequency limit
Figure 9: The conversion of frequencies to wavenumbers are shown for a range of $\alpha$ values. The horizontal dash-dotted line indicates the critical wavenumber corresponding to the path length of the sonic anemometer.
The frequency to wavenumber conversion is a non-linear operation when $\alpha$ is non-zero (see Eq. 9). To assess the impact of $\alpha$ on frequency ($f$) to wavenumber ($\kappa$) conversion, we plot the numerical solutions of Eq. 9 at different $\alpha$ values for a specific 30-min period (Fig. 9). As shown in Fig. 9, at smaller $\alpha$ values, the $f$-$\kappa$ relationship is mostly linear, albeit, when compared to $\alpha=0$, the $\kappa$ values encompass a larger range. The non-linear effects become prominent on $f$-$\kappa$ relationship at $\alpha\geq 0.5$. Apparently, at $\alpha=0.8$, the wavenumbers estimated from Eq. 9 can be as large as $10^{6}\rm\ m^{-1}$ when converting the EC sampling frequency (10 Hz) to wavenumber (Fig. 9a). This cannot be true as it suggests that the EC systems can measure eddies comparable to the dissipation scales, which are clearly smaller than the path length (distance between the transducers) of the sonic anemometers themselves. For our sonic anemometer system, the path length is around 16 cm, and hence the wavenumbers that can be realistically resolved is around $2\pi/0.3\rm\ m^{-1}$. To avoid the aliasing effect, the minimum wavelength being resolved is assumed to be twice the path length (i.e. 30 cm). This limit is shown as a horizontal dash-dotted line in Fig. 9. If we call this minimum wavelength being set by the physical dimension of the sonic anemometer as $\kappa_{\rm cr}$, then from Eq. 9, one can convert $\kappa_{\rm cr}$ to a frequency limit ($f_{\rm cr}$) in the temporal space. Accordingly, one can estimate $f_{\rm cr}$ as,
$f_{\rm cr}=\frac{\kappa_{\rm cr}U_{C}}{2\pi{(1+\kappa_{\rm cr}z)}^{\alpha}}.$ (28)
For our set up, $z=3.1$ m, $\kappa_{\rm cr}=2\pi/0.3\rm\ m^{-1}$, and $U_{C}$ can be found from the space-time correlation curves (see Fig. 4). For the along-wind cases, $U_{C}$ values are averaged over all the five sections. By replacing all these values in Eq. 28, at $\alpha=$ 0.8, $f_{\rm cr}$ appears to be $\approx$ 0.3 Hz (for both along- and perpendicular-wind periods), with some run-to-run variations since $U_{C}$ values are not fixed.
If we consider the $U_{C}(\kappa z)$ values and therefore the derived wavenumbers from Eq. 9 to be correct, then it is unclear what the EC observations represent at frequencies above $f_{\rm cr}$ (i.e. $f\geq f_{\rm cr}$). This is because, beyond $f_{\rm cr}$, the eddy length scales derived from the frequencies are smaller than the physical dimension of the instrument itself. Therefore, it is difficult to associate any spatial scale with frequencies larger than $f_{\rm cr}$. From a practical standpoint, since all our results are from a single height, we cannot establish whether $z$ is the ideal scaling variable to describe $U_{c}(\kappa)$. If $z$ indeed is the correct scaling parameter, then depending on the observation height, $f_{\rm cr}$ values will change (see Eq. 28). We present our conclusions in the next section.
4 Conclusion
In this study, we have developed a comprehensive framework to test the validity of Taylor’s hypothesis and its impact on turbulence statistics over a highly heterogeneous forest clearcut surface. Particularly, our flow situation is symptomatic of a heterogeneous roughness sublayer flow, rather than a canonical atmospheric surface layer one. To execute our objectives, an extensive dataset encompassing both Distributed Temperature Sensing (DTS) and Eddy Covariance (EC) measurements were used. These datasets were collected during a five-month field campaign (May-September 2024) and were separated into two different categories. First, when the mean wind speed was parallel to the forest edge, and second, when it was perpendicular to the edge. For both of these wind scenarios, the conditions were highly turbulent with turbulence intensities of the streamwise wind fluctuations exceeding 0.5. For the parallel-wind conditions, by invoking the elliptic approximation model (he2010small; everard2021sweeping), we showed that the sweeping effects explain why the bulk convective speeds appear larger than the mean wind speed in such highly turbulent conditions. Corresponding to the perpendicular-wind scenario, we could not compute the parameters of the elliptic approximation model with sufficient statistical accuracy. This finding suggests that the accuracy of an elliptic approximation model to describe the space-time correlation curves is impacted in situations when the presence of an upstream forest canopy disturbs the flow over a clearcut location.
Nevertheless, irrespective of the wind direction, we demonstrated that the convective speeds of temperature structures at inertial subrange scales behave in a power-law fashion with increasing wavenumbers. These wavenumbers were directly computed from the DTS spatial data. Contrary to a previous study over a homogeneous atmospheric surface layer flow (cheng2017failure), the exponent of this power law was 0.8-0.8 instead of 1/3-1/3. The 1/3-1/3 exponent was theoretically derived under the assumption of homogeneous and isotropic turbulence at smaller scales of the flow. For the present case in hand, we hypothesize that there could be two reasons behind this disagreement. First, the turbulence remains anisotropic even at smaller scales of flow (chowdhuri2024quantifying). Second, as opposed to cheng2017failure, the conditions that we encounter in a forest clearcut are highly turbulent ($\sigma_{u}/U>0.5$), thereby suggesting that significant sweeping effects associated with large eddies can persist in the flow (everard2021sweeping). These sweeping effects cause modifications to the convection speeds of smaller-scale eddies and one perhaps needs larger exponents in the scale-dependent convection speeds to account for that. In fact, chowdhuri2024quantifying have shown that small-scale anisotropy depends on horizontal velocity variances (see their Fig. 6e), and therefore it is possible that its effect increases with $\sigma_{u}/U$, thereby causing deviations from the 1/3-1/3 scaling. Despite being a promising argument, it is still too early to recommend if 0.8-0.8 should replace 1/3-1/3 in highly turbulent conditions, since more field experiments and large eddy simulation studies over different surface and stability conditions are needed for further verification.
We evaluated the impact of this revised scale-dependent convection speed on the turbulence spectra of velocity and temperature, estimated from the temporal EC measurements. Instead of restricting ourselves to a specific exponent in scale-dependent convection speeds ($\alpha$), we carried out the spectral conversion on EC dataset for a range of $\alpha$ values. This was done to systematically investigate the impact of scale-dependent conversion on the EC measurements of turbulence statistics. Any non-zero $\alpha$ values yield a spectral slope of $\beta+(\beta-1)\alpha$ in the inertial subrange scales (obtained in the frequency domain), where $\beta$ is the spectral slope associated with $\alpha=0$ (i.e. no scale-dependent effects). This finding confirms the assessment of cheng2017failure that applying Taylor’s hypothesis systematically underestimates the turbulence frequency spectrum in the inertial subrange. For our flow conditions, $\beta=5/3$, and therefore when the temporal spectra are converted at an $\alpha$ value of 0.8, they show a broad region of $5/3+2/3\times 0.8\approx-1.13$ spectral slope. This spectral slope enhancement is confirmed from the spatial DTS temperature spectra. To offer an explanation behind this spectral slope enhancement at inertial subrange scales, we conduct a dimensional analysis by considering the influence of a large-eddy velocity scale, namely ${\rm TKE}^{1/2}$, on the inertial subrange statistics.
For the heat flux cospectra, the same conversion yields a slope of $7/3+4/3\times 0.8\approx-1.3$. The appearance of a slope larger than 7/3-7/3 indicates an enhancement of turbulent fluxes at smaller scales of the flow. It remains a topic for future research to evaluate the impact of such enhancement on turbulent flux measurements. However, the non-linear conversion of frequencies to wavenumbers through scale-dependent convective speeds generates unrealistically large wavenumber values. The physical dimension of the sonic anemometers sets a strict constraint on the wavenumbers that can be resolved by the EC systems. In a temporal domain, this eventually results in a critical frequency limit beyond which the eddy length scales derived from the frequencies are smaller than the physical dimension of the instrument itself. This critical limit depends on the power-law exponent of the scale-dependent convective speeds. The existence of this limit dictates that there is a lack of space-time equivalence beyond this critical value.
At present, these results are representative of buoyancy-driven roughness sublayer flows in a heterogeneous location and are also limited by a single observation height. One another drawback is, the estimations of the convective speeds are based on an imperfect temperature signal and we inherently assume that the same could be extended to other scalars and velocity components. It is not immediately apparent whether the structures in all these different turbulent variables move at same speeds. Therefore, for future studies, more extensive spatial data are needed to verify our framework, which ideally should span multiple heights, stability conditions (stable, neutral, and convective), surface roughness (forest canopies, grasslands, arid deserts, water-bodies, etc.), and signal types (velocity and scalars). To achieve this ambitious goal, the observations alone will not suffice and large eddy simulations would be needed to complement the DTS measurements.
Acknowledgements.
[…]
Figure 10: For the along-wind half hour periods from the EC clearcut measurements, (a) the PDFs of velocity and temperature fluctuations are shown. The green dash-dotted line indicates the Gaussian distribution, while the pink dash-dotted line represents the Gram-Charlier distribution. (b) The JPDFs between the vertical velocity and temperature fluctuations are shown where the colorbar represents the contour levels. (c) The flux-weighted JPDFs between the same two signals are shown. (d) The scatter plot between the correlation coefficients $R_{wT}$ and $R_{wu}$ are shown for both along and perpendicular wind cases with respect to the forest edge. The absolute values of $R_{wu}$ ($|R_{wu}|$) are used and the dash-dotted line indicates the 1:1 relationship.
In this appendix, we report the statistics of velocity and temperature fluctuations from the EC measurements corresponding to the along-wind conditions. The conclusions remain the same for perpendicular-wind conditions and hence not being shown. These statistics are used to assess the stability conditions. In practice, the Obukhov similarity parameter is commonly used to determine the atmospheric stability, but we deliberately refrained from that approach since this theory was developed for homogeneous flows and its validity is questionable over heterogeneous surfaces such as a forest clearcut. In Fig. 10a, we show the PDFs of velocity and temperature fluctuations. Typical to a buoyancy-driven flow, the PDFs of temperature fluctuations remain strongly non-Gaussian as opposed to the velocity ones ($u^{\prime}$ and $w^{\prime}$). In fact, the PDFs of temperature fluctuations can be represented through a cumulant expansion method, otherwise known as Gram-Charlier distribution (chowdhuri2019evaluation). This distribution can be expressed as:
$P(T^{\prime}/\sigma_{T})=\frac{1}{\sqrt{2\pi}}\exp[-{(\frac{T^{\prime}}{\sigma_{T}})}^{2}][1+\frac{\mathcal{S}(T)}{6}(\mathcal{S}(T)-3\frac{T^{\prime}}{\sigma_{T}})],$ (29)
where $\mathcal{S}(T)$ is the skewness of temperature fluctuations. The non-Gaussianity in temperature statistics is also reflected in how the JPDFs between the vertical velocity and temperature fluctuations behave. From Fig. 10b, it is apparent that the JPDFs are strongly skewed towards quadrant III, which represents the downdraft motions. To further investigate the heat transport characteristics for our selected runs, we computed the flux-weighted JPDFs, expressed as, $w^{\prime}T^{\prime}P(w^{\prime},T^{\prime})$, where $P(w^{\prime},T^{\prime})$ is the JPDF between $w^{\prime}$ and $T^{\prime}$. The contours of this quantity are plotted in Fig. 10c, and the area under them is proportional to the amount of heat being transported (nakagawa1977prediction). It is immediately clear from Fig. 10c that the transport of heat is strongly dominated by the down-gradient quadrants (I and III) with very small amounts being transported in counter-gradient directions (II and IV).
[…]
Figure 11: (a) The estimations of scale-dependent convection speeds are carried out for the randomly-shuffled DTS data. In accordance with Eq. 5, for the randomly-shuffled DTS data, scale-dependent convection speeds follow a 1-1 slope (shown as blue dash-dotted line). (b) The histogram ($N(\alpha)$) of the scale-dependent convection speed exponents ($\alpha$) is shown for 100 random bootstrapped samples with replacement, corresponding to the DTS temperature signal from Section 6. The vertical line indicates the exponent 0.8-0.8.
For our DTS set up, three types of noises could have impact on the exponent of the scale-dependent convection speeds. The first scenario is a pure white noise. peltola2021suitability showed that at smaller scales of the flow, the DTS measurements could be affected by the presence of a white noise. To test that part, we carried out a randomly-shuffled exercise on the DTS temperature data. At each instant of time, we randomly shuffled the spatial DTS data, thereby destroying any coherence between them. For randomly-shuffled DTS data, the phase differences between the spatial temperature signals at time instants $t$ and $t+\Delta t$ would be equal to $\pi/2$ at all the wavenumbers since no coherence is expected between them. For such scenarios, Eq. 5 suggests that the $U_{C}(\kappa)$ values would vary as ${(\kappa)}^{-1}$, since both $\varphi$ and $\Delta t$ remain constant irrespective of different $\kappa$. This is indeed the case as shown in Fig. 11a. As opposed to the real signals, for the randomly-shuffled samples, one can clearly see that the $U_{C}(\kappa z)$ values vary as ${(\kappa z)}^{-1}$ instead of ${(\kappa z)}^{-0.8}$. However, since the 1-1 exponent is not very different from 0.8-0.8, we carried out an additional statistical significance test. Corresponding to every section of the DTS cable, we generated 100 random bootstrapped samples with replacement of the original temperature data. Subsequently, for each of these bootstrapped samples, to compute the power-law exponent ($\alpha$), a linear regression was carried out between $\log[U_{C}(\kappa z)/U_{C}]$ and $\log[\kappa z]$ at wavenumbers $\kappa z>1$. To illustrate the outcome, in Fig. 11b, we show a histogram of these exponents ($N(\alpha)$) for section-6 of the DTS cable (i.e. perpendicular-wind direction). It is clear that these $\alpha$ values cluster around 1-1 with 0.8-0.8 being located at the far-right-end of the tail (see the vertical dash-dotted line in Fig. 11b). Same conclusion can be drawn for the along-wind sections as well (not shown). Therefore, the exponent 0.8-0.8 is statistically significant with respect to a pure white noise. This is likely since our analyses focused on periods with high temperature variance and hence on periods with high signal-to-noise ratio.
Figure 12: The ensemble-averaged DTS temperature spectra are shown for (a) temporal and (b) spatial domains, respectively. The error bars indicate one standard deviation with respect to the ensemble mean. Only along-wind DTS cable sections (sections 1-5) are used for illustration purposes. The red and black dash dotted lines in (a) indicate the 1-1 and 5/3-5/3 spectral slopes, respectively. The magenta dash-dotted line in (b) indicates a specific slope of $\approx-1.13$ (see Eq. 19).
The second type of noise is related to spectral features of DTS data (cheng2017failure). For the along-wind DTS cable sections, in Fig. 12, we show the spectra of temperature fluctuations computed over the temporal (Fig. 12a) and spatial domains (Fig. 12b), respectively. The error bars indicate one standard deviation with respect to the ensemble mean and the varying shades of gray represent different sections. The spectral amplitudes ($S_{TT}$) are normalized by the temperature variances ($\sigma^{2}_{T}$) estimated from the EC measurements and the frequencies are scaled with the standard surface layer scaling ($2\pi fz/U_{C}$) of kaimal1972spectral. From Fig. 12a, one could see that the DTS temperature spectra in the temporal domain follow a 5/3-5/3 slope at smaller scales while displaying a 1-1 slope at larger scales. Upon comparison of DTS temporal spectra with the along-wind EC temperature spectra, a satisfactory agreement is found between the two (see Fig. S5 in the Supplementary Material). However, since the temporal resolution of DTS data ($2.3\ \rm s$) is lower than the EC one ($0.1\ \rm s$), it does not fully capture the location of the spectral peak. On the other hand, for the spatial domain, the DTS temperature spectra show a rather narrow region of 5/3-5/3 slope (Fig. 12b). This is followed by a region with a spectral slope larger than 5/3-5/3. Precisely, this enhancement is equal to $5/3+(5/3-1)\times 0.8\approx-1.13$ (see Eq. 19).
By following cheng2017failure, we decided to truncate this enhanced region from the spatial DTS spectrum by applying a smoothing filter and reevaluating the phase-dependent convection speeds. This decision was undertaken because upon a closer look at Fig. 5b, it appears that the presence of a 1/3-1/3 slope could be there up to the wavenumber range $\kappa z<3$. We wanted to confirm whether the scaling of scale-dependent convection speeds would revert to ${(\kappa z)}^{-1/3}$ by removing this enhanced region from the spatial DTS spectrum. From Fig. 12b, one could see that the enhancement in spectral powers occurs at scaled wavenumbers of $\kappa z\geq 3$. The limit $\kappa z=3$ is nearly equivalent to a wavelength of 6 m, which corresponds to a spatial separation of 3 m. Therefore, to truncate the spectra beyond $\kappa z\geq 3$, one could average 10 consecutive spatial points, which artificially would reduce the spatial resolution from 30 cm to 3 m. For this artificially reduced data, we computed the spatial DTS spectrum and as expected, the spectrum of this smoothed dataset only revealed the 5/3-5/3 spectral slope (see Fig. S6 in the Supplementary Material). Next, we calculated the scale-dependent convection speeds based on this filtered data to see if indeed those would revert to the 1/3-1/3 scaling of cheng2017failure. This does not appear to be the case, as even for this filtered data, the visual inspection of Fig. S7 in the Supplementary Material suggests that the convection speeds have a preference towards a 0.8-0.8 scaling for all the five cable sections. To confirm this visual assessment, corresponding to the wavenumber range $\kappa z\geq 1$, the exponent was computed by doing a linear regression between $\log[U_{C}(\kappa z)/U_{C}]$ and $\log[\kappa z]$. The answer turned out to be nearly 0.8-0.8 accompanied by an $R^{2}$ value larger than 0.9.
The third type of noise could arise due to a linear-interpolation operation to remove the temperature values at specific locations of the wooden structures. Though the number of points being removed were only 6% of the total cable length, we artificially increased the missing data to be as large as 20% of the cable length. Thereafter, we replaced the missing values with linear interpolation and computed the $U_{C}(\kappa z)$ curves. As shown in Fig. S8 of the Supplementary Material, the results appeared to be insensitive to the linear-interpolation operation. Specifically, with respect to the original signal, one could not detect any appreciable change in the estimations of $U_{C}(\kappa z)$ and therefore, the 0.8-0.8 scaling survived.